A Complete Bibliography of Publications in SIGCSE Bulletin: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

21 April 2018 Version 1.07

Title word cross-reference

3 [LHL79].

0 [NFB^{+}78].

1 [ES78]. 11 [TL79]. 1971 [Jor72]. 1972 [LR73]. 1978 [Wil78]. 1979 [Dal79].

4-year [JTF78].

’68 [CAO^{+}76, Eng76, Con73, EAAD79, LT74, Lit74, Bat74, DN76, GG73, Men74].

’70 [Zin70]. ’77 [WP78b, Lop78]. ’78 [EAAD79, DAB^{+}79].

a.k.a [Eck78]. ability [New75, Tro79]. Abstract [Bol73, Eng76, Kug75].

Abstracts [Zin70, Bun71]. Academic [MS76b, Bar75b, CD79, KL78, Miz76a, RH73, TB77]. academic/industry [CD79]. Academy [Wu74]. accelerated [Dal77]. acceptable [UA76]. achieve [Nei77]. ACM [Ash73, ABB^{+}77b, Ban78, Eng71a, Hug70, LSL^{+}76, LSA^{+}76, Lop78, SLC^{+}74, Sap75, Wor78]. Activity [LSL^{+}76]. Adapted [Dal76]. adaptive [Sam74]. addition [Mod74]. addressing [Bei72]. Adequate [Hug70, Wes72]. Adleman [BB79, SP79]. administrators [God75]. admissions [CMT74]. adopts [Tho75]. Advanced [Kha77a, Ham75a, Lov72, Lus77, McF75, TT78b, Wil76a, vDSMM74]. advising [Gil76]. Advisory [GH79]. aerodynamics [For73]. aid [BBS75, MD76, PP77, RJB71, Wol76a]. aided [KF76, Mar72, SS79]. aiding [Maz79]. AIDS [Cha78c]. algebra [SS79]. ALGOL [KR72]. ALGOL-E [KR72]. ALGOL68 [Hed77]. algorithm [Jon79, Kon75, Nor74, Pum75]. Algorithmic [Ott76, IE72]. algorithms [Ano77, Ker78]. Alice [Car78]. alone [Bow78]. alphabet [Mod74]. Alternative [Kha77b, New73, DN76]. alternatives [Rin78]. always [BB79]. American [Mij78, Rot78]. analog [KPM74]. analog/hybrid [KPM74]. analysis [Cor78, EDL70, Hed72, IN76, IE72, Ort74, Rot75, SG78, Whi72, YGM73]. Animated [Bae75]. annual [Ing75, McF75]. APL [KH74, PW74]. application [Bai73, Pow79, SI79]. application-oriented [SI79]. Applications [Coo77a, BL76, Cab73, God75, KW78, Lak76, Lov79, PW75b, SBS76, Sch78b, TM74, TS75]. applied [Lus77, MF79]. Approach [Ott76, Aik78b, AHP77, Beu75, Che76c, Con77, Coo77b, CL76, Dal78d, Gol71, HH73, Jär77, KR72, Kim78, Lan75, MF79, MH75, Mit79, Nel78, PW75a, PB77, PS78c, Pum75, Rag79, Seb76, SM76, SS74, SW79, UA76, Vis77, Yel78, Aik78a]. Approaches [AT77a, Lem79]. Arabia [Air76]. architecture [Slo74, Smi76, Tho74]. area [Wes72]. areas [Oli74]. arithmetic [Mat70b]. arrogant [DS72]. Articulation [JBF76]. artificial [Fin70]. arts [ASW78, BNS78, Din75b, Kom70, LR72, LR73, Lei72, Mei79, Nan76, Rot72, Sap75, Smi78, Smi79, Wor78, Zim71, dC70]. ASAG [WA74]. ASAP [MK73]. Assembler [PP77, Bas77, Bei73, Bei77, Mas73, McC74, Sjo76]. assembly [Smi76, vDDDN76]. assessment [Bra70, Fre76a]. assignment [Mac79, WA74]. assignments [Aik71a, Wei73]. ASSIST [HP76, Mas73]. ASSIST-V [HP76]. assisted [IN76, Lew74, SBS76, Swi76]. Association [LAS^{+}77]. asynchronous [Vis78]. Atlanta [MS76b]. ATOPSS [Tra74]. attitude [Luc76]. attitudes [Bai73, PW74, Slo76, Wai79, Wil73]. audio [SS79]. audio-tutorial [SS79]. augmented [Gro75]. Australia [Moc78]. author [WA74]. author-mode [WA74]. authors [Shn77]. automata [Aik70, RJB71]. Automatic [Wol76a, DN75, MM73, Mus76, NW74]. Automating [NR73]. awareness [Maz77].

B.S [Kal74]. B.S. [Map73]. B3 [Kor75, Wax75]. B4Tran [MH75]. Bachelor [Arc75, KK78, SW71]. Background [Hug70]. Baker [PS78b]. base [GH78, Lam78, LT78, Ste78]. based [AT77a, BR79, Bow78, FL76, GG77, Sap75, Vaw76, Wol76b]. basic [BB76a, Beu75, Coo78, Rag79]. Basic2 [Dal78b]. Battelle [BJ76]. battered [Har77]. Bayesian [IN76]. be [HM75]. BECUN [BJ76]. before [Kim79]. beginners [NS76, Tay77]. beginning [Bre76, Kei76, Rot78, SW74a]. Behavioral [SI79, Bal78]. belong [Din75b]. between [Ega76, TT78a]. binary [Den79b]. black [MS76b]. blind [Har77]. board [Lev79]. book [Bar75a]. both [NS76]. BRAINO [AS78]. Breadth [SSL79]. bridge [Web75]. Brigham [Hay74]. Buffalo [Fin70]. Bulletin [Lit77]. Business [AP70, Cha74a, CS77, DEN79a, Hon75, KH75, Miz76a, Sch78b, Ste78, Fai74b]. Business-Oriented [Fai74b].

C [LHL79]. CADA [IN76]. CAI [DFF^{+}78, Lin76, LK76, Sob76, Wu74]. CAI/CMI [Lin76]. CAJC [LHL79]. Calculator [Mou77]. Calculators [Mou76]. Calumet [AM79]. campus [AM79, Mod73]. can [Shu79]. Canadian [Slo74]. can’t [LL78]. captioned [MPS78]. card [Der70]. Career [BCS79, LSL^{+}76, LAS^{+}77]. Careers [Haa78]. case [Car78, Chv78, Ham75a, Hom77, Kor75]. cases [Lov79]. center [Pec77, Var78, MS76b]. central [Eck76a, MHA78]. certification [BFC79, Har74]. challenge [DKMS75, HWC75, Moo70, Smi79]. change [Luc76]. changing [NW74]. chemical [Mor72]. Chief [PW75a]. Child [Sol76]. City [Aik78a, May75]. clarity [Den79b]. class [Kei76, PW75a]. classes [KS78]. Classroom [Bau76, Bau75, KR76, McG73, Mil78, Wel76]. Closing [Ega76]. CMI [Lin76]. CMU [Dep73]. COBOL [FM78, Kei76, Kha76, Kha77a, Miz76b, BC77, Lem78a]. College [Aik78b, CMT74, LR73, LHL79, LSL^{+}76, LSA^{+}76, Con73, DRT75, Din75a, Din75b, Ell73, HT73, Lei72, LAS^{+}77, Mor77, Mou74, RH73, Rot72, Sap75, Smi79, Sti73, SS79, Tho75, War75b, Wol76b, Wor78, Zim71, dC70, GM79, May75]. colleges [ASW78, AE72, CK79, Con75, Eng70b, Eng71b, JBF76, LR72, LR73, Lew75, Lew76, Lit74, Pec75, PDS^{+}79, Weg71]. Combining [BS76]. COMES [DFF^{+}78]. Comments [Eng70a, Hos73]. commercial [Sob76]. Committee [Eng71a, LHL79, LSA^{+}76, MK74, ABB^{+}77b, Cai75, Ham79, LAS^{+}77]. Communication [San78, San78]. communications [Thu79]. Community [CMT74, LSL^{+}76, LSA^{+}76, BR79, Con73, JBF76, Lit74, LAS^{+}77, LHL79]. community/junior [Con73]. comparative [Pag78]. Competency [KRFS79]. Compiler [App79, Cal78, Cha75, Bar75a, Dal78d, Din75a, FS79, HHP77, SM76, vdR79]. compiler.kit [Abb77a]. Compiling [TT78a]. complementary [Kim78]. complete [BC70]. component [Der78, Mat70b]. composition [Kim79]. comprehensive [Cla79, Cor70]. computability [Har78]. computation [CC72]. computational [Mat70b, Rhe73, YGM73]. Computer [Abb75a, Agr76, Aik78b, Air76, Any78, Atc73, AE72, Bae75, BC79, BP73, BM75, Boy76, BNS78, Cab73, CK79, Cha74a, CMT74, Chv78, CH78, Cur75, Dal79, EG76, Eng70b, Eng71b, Eng71a, Fle74, FGH^{+}70, FK73, God74, God75, GG73, Haa78, Ham75b, Ham77, HBJ76, JRS78, Jor72, Kal74, Kay74, LR72, LR73, LT75, LHL79, Lei72, Lev76, Lew76, LSL^{+}76, Lov75, Mal72, Mar72, MD76, Mei79, Mod73, Mor72, MK74, MN76, Mul76, Nah79, Nan76, Nei76, NT70, PE73, PB74, Poi79, Ran75, Roo72, Shn71, SS74, Slo74, Slo76, Sny70, Sol76, ACM72, Sta72, SAK^{+}78, Sto74a, SS79, Tho74, Var79a, Wal73, War75b, WP78a, Abb75b, Abs78, AH72, AIS72, Aik71b, Aik72, AC72, And72, Ano76, Arc75, AP70, Ash73, Aus74, ACEH76, ABB^{+}77b, Bai73, Bar75b]. computer [BMOW76, Bat73, BJ76, Bei74, BG76, Ber79, Beu75, Bon74, BL76, BC70, BS76, BFC79, CJPT77, Car78, CC72, Cha77, Cha78b, Che75, Che76a, Che76b, Cla79, Cla74, Cla76, Con77, CAO^{+}76, Coo77b, Cor78, CS77, CR70, Dal77, Dav77, DP79, Den70, Dep73, Der78, Des72, DAC^{+}78, Din75b, DG74, Dow74, DMH^{+}77, Eck72, Ega76, EDL70, EW72, Eng70a, EB72, EJ73, Eps75, Fai74a, FAT78, FL76, FHSW78, For70a, FJ78, FJ77, Gib77, GLO74, God73, Gol71, GS76, Gon78, GB73, Gra78, Gro75, Ham74, Ham75a, Ham79, HLEA79, Ham70, HAB^{+}76, Ham75c, Han78, HM75, Har74, Hay74, HA71, HJ73, Hel73, Hof74, Hom78, Hon73, Hon75, Hos73, Hun70, IN76, Iye76, Iye78, Jac76a, Jeh74, JBF76, Jon75a, Jon75b, JK77, KK78, Ker78, Kha77b, KL78, KH75, Kim78]. computer [KRFS79, KF76, KH74, KS74, Kru78, KPM74, Lak76, Lav78, Lev79, Lew75, LI76, Lin78, Lin79, LAS^{+}77, Lov79, Lov74, Lov72, Luc76, Lus77, MF79, MM78a, MS77, Man74, MK72, Mat78, Mat70b, Mau71, Maz76, Maz77, Maz79, MB76a, MB79, Mil75, Mil78, MM78b, Moc78, Mod74, Moo70, MHA78, Mou76, Mou78, MPS78, Nei77, New73, Nie70a, Nie72, Nie70b, NR73, NFB^{+}78, Noe73, Noo79, OR77, Oli73, Oli74, Oml78, Pag78, Pee76, PS78a, PB73, PB77, PC76, Poi76, Pol78, PK74, Pos79, Pow79, Rag79, RH73, Ral71, Rei71, Rin78, RGH78, Rob75, Rog78, Ros76, Rot72, Rub76, Sal77, Sam74, SBS76, San78, Sch78a, Sch79, Sch77, Sch71, Sch73, Sem73, Sen74, SI79, Sha72, Sha73, Sjo76, SR76, Smi79, Smo76b, Sol79, Som77, Son75, Son78]. computer [SG78, Ste72, SP70, Sti73, Sto78, SSL79, Sto75, SW71, Swi76, SW79, Tch75, TT78b, Tho79, Tho75, Thu79, Tra76, TM74, Vaw76, Vic71a, Vis77, Wai79, Wal74, War75a, Wea78, Wea79, WZ73, Weg70, Weg71, WW74, Wei77, Win78, Win77, Wol76b, Wor78, YGM73, Yel78, Zel72, dC70, Lin76, AM79, CM75, GH79, Pec75]. Computer-aided [SS79, KF76]. computer-assisted [IN76]. computer-based [FL76, Wol76b]. Computer-controlled [Lev76]. computer-graphics [Dow74]. Computer-scored [Boy76]. computer-supported [Jon75a]. COMPUTER-TUTOR [Lin76]. Computers [ACEH76, Con75, HH73, Kol79, Lee72, Mou74, Mou79, Poi78, PDS^{+}79, Sam75, AP71, Bol73, Dal76, For73, GG78, Iye78, Kom70, Mar76, McG73, Mor77, Rhe73, Shu79, Wil73, Zin70]. Computing [BKDH79, Hob78, LAS^{+}77, Wil76b, Zim71, ASW78, Bra70, CGS75, DKMS75, Eck76a, Fre76a, HWC75, Ing75, Lem78b, Lew76, May75, Mil73, MS76b, Mul76, Nie70a, Sap75, Smi78, Sol78, SW74b, TP72, Tay79, Tha79, Var78, deG72, Fai74b]. conceal [BB79]. concentrated [Ell73]. concentration [Oli74]. concept [Con77]. concept-oriented [Con77]. concepts [Aik72, Eck76b, Lam76, Lev76, Miz76b, Tra74, Wak78]. concrete [Kug75]. concurrent [Dal78c]. Conducting [KS78]. configuration [Rei71]. confusions [Ham75c]. connection [TT78a]. Consequences [DAB^{+}79, Noo79]. Consideration [Fre76a]. considerations [FK76]. considered [DS72, GCS77]. consortium [MS76b, BKDH79]. constrained [Fre76b]. construction [Cha75, Dal78c, SM76]. contain [HFL^{+}78]. content [God73]. Contesting [Met79]. continued [For70a]. continuing [Des72]. control [EB79, KPM74]. controlled [Lev76, SSL79]. Cooperative [AH78, Arn78, HA78, FJ78, God74, War75a]. core [Cla74]. Cornell [Sal73]. Correction [ACM72]. correctness [Mau77, Yel78]. cost [Lem78a, UA76]. cost-effectiveness [Lem78a]. Council [GH79]. counterfeit [WW74]. countries [Chv78, Maz79, San78, SAK^{+}78]. country [Lav78]. Course [Hof74, Mat70a, Nar70, Aik70, Aik71a, Bai73, Bar75a, Bas77, Bau79, Bei73, Bei79, BL76, Bow78, Cal78, Cav75, Cha75, Cha77, Cha78a, Cha78b, Che75, Cla76, Con75, Coo76, Coo77b, Coo77a, Dal76, Der70, DRT75, Din74, Din75a, Din77, Dow74, Dun74, EB79, ES78, FHSW78, FM78, FM79, Fre76b, FJ77, Gib77, Gra78, Gri74, GG77, GG78, HM75, HSW78, Hom77, HGU79, Iye76, Iye78, Jal78, JW78, Ker78, Kha76, Kha77b, KH75, KW78, Kol79, Law74, Lee72, Lev78, Lin76, Lov72, Lus77, Mac74, Mar76, MT73, MM73, Mat78, McC74, MB79, Miz76b, Moo79, Nan76, New75, Noo79, Pag78, Par72, Phi76, Poi76, Poi78, Pra76, PS78c, Pur75, Rad76, Ral71, Rip75, Rus77, Sam71, Sam75, Sam74, Sch78a, Sch78b, SM76, Sha72, Sha73, Smi78, Smi76]. course [SB74, SP70, Ste78, Sto78, Sto74b, SW79, Tha79, Thu79, TM74, TS75, UA76, Vic71b, Wak78, WW74, Wei73, Wil76a, Win78, Win77, Wol76b, Yel78, vdR79]. courses [AH72, ASW78, AHP77, ACEH76, Bra70, CGS75, Coo78, Ell73, Gin75, God73, Gri78, Han78, Hof78, JK77, Lov74, Men74, Mit79, Mou78, NR73, Rhe73, Rog78, Ros76, Sol78, SSL79, Sto74a, TT78a, TT78b, Wil76b, vDSMM74]. coursesr [Abb75b]. creative [Kay74]. critical [Bra70]. critique [CAO^{+}76]. Crosse [WP78b]. cryptosystem [SP79]. cryptosystems [BB79]. CS1 [Bow78]. CSA [Wil78]. culminated [Sch78b]. Culture [Sol76, Ros71]. CUPM [Rhe73]. currency [CH78]. Current [JRS78, Con75]. Curricula [Fai78, Kea75, Rin78, DAC^{+}78, JBF76, LHL79, Oli73, Pow78a, Rub76, Sap75, Vic71a, Wal73, Wal74]. curriculua [CK79]. Curriculum [ABB^{+}77b, Eng71a, Eng74, Eng76, LSL^{+}76, LAS^{+}77, Lop78, SB74, SHLS79, AH72, AIS72, AM79, Ash73, Bal78, Ban78, Bar75b, BG76, Cla74, CAO^{+}76, Cre78, Dal78a, DP79, Eng70a, EB72, EJ73, Fai74b, Ham79, HAB^{+}76, HHD78, HBJ76, HA71, HT73, Hof79, JTF78, KH74, Lam78, Lev79, Lin78, LSA^{+}76, MS77, Mat70b, MB76b, Mil75, MM78b, Miz76a, Mou76, Pag78, Pec75, Rot72, Rot78, Sch71, SR76, Som77, SG78, SW74b, Tho74, Tho75, Var79a, Wea78, Wor78, Bat74, CM75, Con73, DAB^{+}79, DN76, EAAD79, GG73, LT74, LHL79, Lit74, LSA^{+}76, Men74, WP78b].

D [Jor72, CC72, Mau71]. D.P. [Hug70]. Dame [Sch71]. dans [San78]. Dartmouth [GM79]. Data [Abb75a, Iye76, LHL79, Vic71a, AT77b, BP77, BR79, Col79, Din74, GH78, Gum77, Hag76, IN76, Kea75, KS78, Lam78, LT78, MF79, Man74, Phi76, Rot78, Smi76, SB74, Ste78, TS75, Wes72, BM78]. data-oriented [Col79]. data-processing [Kea75]. Database [Dal78a, Pur75]. Databases [TT78a]. date [FAT78]. Dayton [Dal79]. deaf [BP77]. debugging [BBS75, HA71, Lem78a, Mat74, PP77]. decision [JW73]. declarative [Ada75]. defined [Nah79]. definitions [Nei77]. Degenerate [Ber09]. Degree [ACM72, Arc75, GM79, Gon78, HLEA79, HAB^{+}76, Jon75b, MM78a, PE73, SW71]. delivery [Fre76a]. demand [Ham75b, Ham77, Lei77]. democratic [Whi72]. demonstrate [For73]. demonstration [KHKD78]. demonstrations [Lev76]. Department [MK74, AP71, AP70, Lav78, Sto74a]. departments [Ash73, Nie70a, Slo74, Vic71a]. departments/curricula [Vic71a]. Depth [Ano77, SSL79]. Depth-first [Ano77]. describing [Jac76b]. Description [Law74, Sam75, Cla76, FM78, FM79, Kol79, Lee72, Lin79, Mar76, McC74, Poi76]. descriptions [Mat70a, Nar70]. Design [Pol78, Wel76, Bau75, Cha74b, Col79, Coo77a, Din77, EW72, Fre76b, Gra70, Gri78, Gyl78, Jon79, Ros76, Sal77, SG78, SS79, Var79b, Vis77, Vis78, Wak78, Whi77, WP78a]. designed [Bau70, Hug77, Mil75, Pal76]. Detection [Ott76]. determining [Nor74]. developed [Maz79]. dévelopement [San78]. Developing [Maz77, Any78, Chv78, Lav78, Rob75, San78, SAK^{+}78, Sto75]. Development [App79, MK72, May75, Sap75, Air76, BR79, Irb77, Pec77, Pum75, SB74, SHLS79, TMWD77, War75a]. devices [JK77]. diagnostics [DTW75]. diagram [KF76]. diagrams [BG74, Smi76]. dialogue [AP70]. differential [For73]. difficulties [Cha75]. digit [Car78]. digit-Alice [Car78]. digital [Cai75, Chu75, SS74, Vis77, WP78a]. digraph [Ano77]. Dijkstra [DS72]. dimensions [JRS78]. Direct [TT78a]. directed [Pow78b]. directions [SLC^{+}74, YGM73]. disc [Mic76]. disciplinary [Rag76]. discipline [FJ78, LK76]. disciplined [CN76, Emb78, Fri75]. disciplines [God73]. Discrete [EJ73, Ber76, DRT75, Pra76, TM74, Wax75, YGM73]. Discussion [Ano76, BFC79, DAB^{+}79, EAAD79, Ham79, HLEA79, Maz79, Met79, Mou79, PDS^{+}79, Bon74, FBB^{+}76, GKC^{+}74, McF75]. disk [MHA78]. display [Che76b, Lev79, Lev76]. Dissertation [Bun71]. distributing [Eck76a]. diversified [Che76a]. division [BG76]. do [Abb77a, BB79, Mou79]. do-it-yourself [Abb77a]. Documentation [Bre76]. Does [Din75b, Maz76]. doing [PDS^{+}79]. down [GG77, Hed72, Mar75, Pum75, Rag79, TB77, Wil72]. dozen [PS78b]. DRAW [Tho79]. drill [LK76]. Dynamic [Sch73, Kay74, Lev76]. dynamically [Whi77]. dynamics [HHD78, Mar72].

Eastern [Hag76]. economics [Phi76]. Editorial [BD74a, BD74b, Bon75]. EDP [Cab73]. EDP-applications [Cab73]. Education [Dal79, LHL79, AH78, Ano76, Any78, Arn78, AE72, Aus74, Bar76, Bol73, Bon74, BNS78, BS76, Cab73, Cai75, Cha74a, Chu75, Chv78, CAO^{+}76, Der78, Des72, DG74, EDL70, Eng70b, Fai74a, FAT78, FJ78, Fre76c, FK73, GS76, GG73, Ham70, Han78, Har74, HA78, Hay74, HJ73, Hos73, Jac76a, Jeh74, JRS78, Kim78, KS76, Kom70, Lei72, Maz79, Moc78, Mod74, Mou74, Mou77, Mou79, Mul76, Nie72, NT70, Noe73, Oli74, Pec77, PS78a, Poi76, Poi79, Pos79, PDS^{+}79, Rob75, SI79, Shn71, SAK^{+}78, TP72, Tho79, Var79a, Weg70, Weg71, Wei73, Wil76b, Zin70, deG72, MK74]. Educational [BP77, BKDH79, Cha78c, BJ76, Gla75, Jär77, Kay74, KP76, Lew76, LT78, Mic76, Mor77, Son78]. educators [Tay79]. EDUSAC [KP76]. Effect [MS77, EB72]. effective [Lev77, Rip75, UA76]. effectiveness [Lem78a, Var78]. efficiency [AT77b]. effort [War75a]. efforts [Zim71]. electrical [AP71, Slo74]. elementary [Fri75, FK76, MN76, Mou77, Tra74]. elimination [SR73]. emergence [Mat70b]. emphasis [MB79]. emphasizes [Bei72]. encryption [Ber09]. enforcement [Poi78]. Engineering [Fai78, AP71, AHP77, Bar75b, Bol73, DS75, Fre76c, Gon78, Hof78, Hof79, Jeh74, JRS78, JTF78, Par72, RH73, Rin78, Slo74, Son75, WZ73]. engineers [BC77, Dow74]. English [Kim77]. Enhancing [GS75]. enough [DP79]. enrollment [BC79]. enseignement [San78]. entry [LHL79, Wei77]. entry-level [Wei77]. environment [CMT74, EB72, FM79, Mei79, Mij78, PP77]. environments [BB76b]. epsilon [Bat73]. equations [For73]. equilibria [Mor72]. error [DTW75, McC78]. essential [Nan76]. established [Pow78a]. Etc [ACM72]. Evaluating [Shn77]. evaluation [Cla76, Fle74, FK79, Hag76, HA71, Noe73, Rad76, Wal74]. even [PW74]. Evergreen [Aik78b]. Evolution [RH73, Sem73]. exam [Fin70]. Examination [Ham75a, Cor70, Nie70b, Whi77]. examinations [For70a]. example [PW74, Pum75]. examples [Aik71a, Seb76]. Exams [Roo72, Aik71a, Ano70a, Ano70b, Ano71, For70b, Lio70, Mau71, McF75, Rei72]. exchange [Son78]. Execution [Bae75, KHKD78]. Exercises [Jor72, Rag76]. Experience [LT74, SP70, Sti73, TL79, Bar75a, BC77, Der78, Fri75, Iye76, Iye78, Mas73, Men74, PS78a, PL78, TB77, Wai79]. Experiences [Aik72, BT74, Bau79, GS76, GG73, GH79, Bra70]. experiment [Cha77, Mav76, SSL79]. Experimental [FK79, AC72, Din74, Hom77, KR72, KR76, PC76, Wil76a]. explanatory [NFB^{+}78]. extended [Gum77]. extensibility [Gum77].

F [Jor72]. faced [Lav78]. facilitate [Hug77, Pal76]. Facilities [Ano76, Con75, MS77, Zim71]. facility [GS76, MHA78, Ran75]. facing [Rob75]. Factionalism [Lem78b]. faculty [BC79, Mor77, Rot75, Wai79]. Failing [Hug70]. features [FK79]. February [Dal79]. federal [Cur75, May75]. Federighi [Jor72]. field [OWE73]. fifth [Ing75, McF75]. File [Iye76, Ric75, MK73, Tha79]. files [Moo79]. Final [MK74]. First [FJ77, Ano77, FAT78, GG78, HM75, HGU79, Mat78, McC74, Moo79, PS78c, Ral71, SM76, Smo76b, SW79, Tho75, TM74]. first-year [Smo76b]. fixed [SP79]. FJCC [Hug70]. flow [KF76, OWE73, Smi76]. flowcharting [Bre77, Hos77]. flowcharts [HBJ76, JW73]. foreign [BP73, Cal71]. formal [Aik70, Har78]. formulation [MB77]. FORTRAN [Bez75, FK76, FK77, Wag75, Coo76, Emb78, Gum77, LT76, Maz78, MH75, NS76, PP77, Rad76, Ral71, Ros71]. forward [Mou74]. Foundation [BMOW76]. Foundations [Wei77, Hon75]. four [HT73]. four- [HT73]. franca [KH74]. Francisco [Haa78]. Fraser [SW74b]. fraternalism [Lem78b]. free [McC78]. French [San78]. freshmen [DS75, Kim77]. function [Kra72]. functional [Nei77, Sal76]. fundamental [FHSW78]. fundamentals [YGM73]. funding [HFL^{+}78]. further [RD77]. future [SLC^{+}74, deG72].

gambits [Bar78]. games [Der70]. gaming [BB76b]. Gap [Ega76, CD79, Web75]. Gauss [SR73]. general [Tra76, UA76]. generalistic [Kim78]. generalized [SB76]. generating [Pal76, Wak78, WA74]. generation [Mus76, Wei73, Wol76a]. geographical [Wes72]. geometric [MD76]. given [Sam75]. goals [Mou77]. good [LM76]. government [CAO^{+}76, Pec77]. GPSS [Lam76, RD77]. Grade [New75]. grading [MM73]. Graduate [Ham75a, Tay79, Aik71b, BCS79, BC77, Cha78b, Fai74b, Gin75, GB73, Iye76, Iye78, Kal74, KK78, LI76, Mat78, Mij78, Pur75, Sch79, Smo76b, Weg70, Wei77]. Graduates [Haa78, Abb75a, Hom78]. graph [Hon73]. graphical [Jac76b, SS79]. graphics [Cor78, Dow74, MD76, Tho79]. GRE [McF75]. Gries [Bar75a]. group [CKP78, FK73, Moc78]. guide [Gin75, Shn77]. guided [Coo77a]. guidelines [DAC^{+}78, LHL79, LAS^{+}77]. guiding [CS77].

hacker [Har77]. hand [AP70]. hands [Aik78b, CL76, Lan75, WW74]. hands-on [CL76, Lan75, WW74]. Hardware [Cla74, Gla75, HFL^{+}78, Fai74a, Hob78, PK74, Pos79, Sal77, WP78a]. harmful [DS72]. health [BL76, Lak76]. heart [GLO74]. held [Ing75, LR73]. help [MM73]. helpful [GCS77]. hidden [Sob76]. hierarchical [GH78]. High [Tho75, Aik72, BM75, FK73]. High-school [Tho75]. Higher [MK74, Sny70, deG72]. hireable [Cab73]. historical [Kim78]. history [Hom77, Rog78]. HMS [WW74]. home [WP78a]. honor [Bat73]. human [DKMS75, HWC75]. Humanities [Sam75]. hybrid [KPM74]. hypothetical [Rei71].

I/O [Bai72]. ideal [BNS78, SP74]. ideas [Kug75]. IFTRAN [Bez75]. II [Ker78]. III [MK74]. IIT [MB76a]. Illinois [LR73, KRFS79]. image [KL78]. immediate [TP72]. Immigration [Sha72, Sha73, Cha78b]. Impact [Bol73, CL76, Rhe73]. Implementation [Ban78, CKP78, HP76, Lop78, Rus77, Sap75, Sto77, Whi77, WP78b]. Implementing [MM78b, SR73, DAC^{+}78]. Implications [Ash73, FK79, FAT78]. improved [CL76, Nel78]. In-service [Jac76a]. including [FAT78]. Increasing [Den79b]. independent [Bei73, Wol76a]. index [Var78]. Indiana [Man74]. Individualized [Gra78, KS74]. individually [Pal76]. Inductive [Dey73]. Industrial [Sch78b, DKMS75, HWC75, Kor70, WZ73]. Industry [DG74, Bon74, CD79, CAO^{+}76, Ega76, Ell73, God74, Kor70, Mil75, Web75]. Industry/University [Web75]. inexpensive [Dal78c]. informal [Abb75b]. Information [Kal74, Nei76, Ash73, BBS75, BS76, HT73, Map73, MK72, Mij78, SB74, SHLS79, Vaw76, Vic71b]. informatique [San78]. Initial [Eng74, Eng71b]. Innovative [Lew75, Sto75]. Input [Eng71a]. instant [Abb77a]. Institute [BP77, GG73]. institution [Ing75, Sto75]. institutions [Ano76, Cab73, Jon75a, MS76b, Sny70]. instruction [Bar78, Che76b, Con74, Eng71b, FHW73, Hed77, Jon75a, KRFS79, KS74, Lew74, Mar72, Pal76, SBS76, Sen74, SB76, Swi76, Wes72]. Instructional [KR76, Bai73, BB76a, DTW75, Eck78, FBB^{+}76, FL76, JK77, Lam78, Lov75, Mil78, Pol78, Sal76, Sch73, WM74]. instructionally [UA76]. instructors [Shn77]. instrumental [TMWD77]. Integrated [SW79, HT73, MHA78]. integration [FJ78, HGU79, Men74]. intelligence [Fin70]. Intelligent [FBB^{+}76]. Interactive [DTW75, SW74a, Tha79, Cor78, FL76, Gil76, Maz78, MS76a, Sjo76]. interdisciplinary [Agr76, Aik78b, HH73, Kru78]. interest [Ing75]. interface [NW74, WA74]. intermediate [TT78b, Win78]. internal [Lev77]. internship [BS76, Pow78b, SCM78]. internships [God74]. interpreter [BB76a, Lev78]. Introducing [Cou75, Gol71, HSW78, Men74, dC70, BC70, Rag79]. Introduction [SS74, Col79, Eck76b, PB77, Wak78]. Introductory [AH72, Che76c, Con74, Ros76, Sal73, AHP77, Bau79, BL76, Bra70, CGS75, Cha77, Che75, Con77, Coo76, Coo77b, DN75, Der70, EB79, ES78, FHW73, FM78, FK79, Gib77, Gri74, Jal78, Kha76, Kha77b, KH75, Kim78, Law74, Miz76b, New75, NR73, OR77, Pow79, Rog78, Sam74, Sch78a, Seb76, Shn77, Sjo76, Smi78, Smi76, Sol78, SP70, SSL79, TT78b, TS75, UA76, Wak78, Wei73, Win77, vDSMM74]. involved [AH78]. IPI [SS74]. IPL [Sch77]. Island [CMT74]. Isolated [Shu79]. ISP [Cou75]. Israel [Sny70]. Iterations [Mil78].

Jackson [Lew76, LI76]. Jersey [MK74]. johny [LL78]. Joint [Jeh74]. Jose [Nei76]. Josephus [Woo78]. Jr [Jor72]. July [LR73]. Junior [LHL79, LSL^{+}76, LSA^{+}76, Con73, LAS^{+}77].

Kansas [Nel71]. keep [Bre77, FAT78]. Kentucky [Hag76]. key [BB79]. keys [Ber09]. kit [Ran75]. KWILT [Eck78].

laboratories [Hob78, HFL^{+}78]. Laboratory [Jor72, Dal78d, Eck72, FS79, GS76, GG77, Hun70, Irb77, Kru78, MS77, MHA78, PS78c, RGH78, Sol79, Sta72, TMWD77]. Language [Gum77, Wel76, Aik70, AS78, Bas77, BT74, BP73, Bei73, Cal71, Cha74b, Dal78b, Dal78c, EB79, Eck78, FHW73, FK79, Har78, IE72, Kha77b, Lec74, McC74, MH75, MK73, Oli74, Sen74, Smi76, Som77, TL79, WM74, vDDDN76, vdR79]. languages [Dey73, DFF^{+}78, Din74, FK79, KR72, Law74, Lem79, Mav76, Pag78, Rei73, Sch73, SR76, Sol78]. large [Che76b, CR70, MM73, Vic71b, WW74]. large-scale [CR70, WW74]. Latin [Mij78]. latter [Lem78b]. law [Kol79, Poi78]. Leading [Sol76, Har77]. Learning [BG74, Pee76, Swi76, HHP77, Sny70, Vaw76]. lecture [PS78c]. lecture/laboratory [PS78c]. lectures [DEN79a]. left [AP70]. left-hand [AP70]. less [Maz79]. lesson [AT76]. lessons [Cor78, SS79]. level [Aik71a, Fri75, Iye76, Iye78, Mou78, Nie70b, PB77, Vis77, Wei77, Wol76b]. Liberal [Kom70, ASW78, BNS78, Din75b, LR72, LR73, Mei79, Nan76, Rot72, Sap75, Smi78, Smi79, Wor78, Zim71, dC70]. libraries [KW78]. Library [Lov75]. limited [War75b]. line [Che76b]. lingua [KH74]. LINUS [WM74]. list [KP76]. literacy [Aik78b, ACEH76, CS77, Luc76, Nei77, Wai79]. literature [Gin75, Gol71]. little [vdR79]. locked [KPM74]. logic [Vis77]. logical [Vis77]. long [Cla79]. long-term [Cla79]. look [Pol72, Pra76]. loops [KPM74]. low [Nie70b]. LSI [Hob78].

M.S [AP71, Ber79, GB73, Hel73]. Machine [Som77, Wei73, AS78, Bei73, Din74, MS76a, NW74, Smi76]. Machinery [LAS^{+}77]. machines [RJB71, SP74]. Macrognosographics [Bre77]. macromodules [EW72]. maintenance [Bei79]. major [Sap75]. majors [Coo77b, Dav77, Gib77, Lov72, Oml78, Sam74]. make [Kug75]. Making [Miz76a, HA78, Whi72]. man [NW74, Pow78b]. man-machine [NW74]. Management [Iye76, BCS79, Dal78a, GH78, LT78, Pur75, Var78, Var79a]. managers [Des72]. manipulation [Fin70]. manpower [Air76, Ham75b, Ham77]. maps [Var79b]. mass [Bar76, Wei73]. Master [MK74, GM79, Hof79, KK78, MM78a, Nei76, Web75]. Masters [Ham79]. materials [Sha73]. mathematical [CC72, Dav77, Lov74, RJB71]. mathematics [Arc75, FGH^{+}70, Gin75, NT70, Pee76, Rhe73, SB76]. Maximal [TL79]. measurement [Cla76, Noe73]. Meditations [Smo76a]. medium [Kay74]. members [Mor77]. merging [MM78b]. messages [BB79, SP79]. metaphors [Mou77]. method [Mil78, Wu74]. Methodology [OR77]. Methods [Ger75, Dey73, Fle74, Wil76a]. Mexico [Gon78]. Microcomputer [Coo78, Lev79, Hom77, Sol79]. Microcomputers [Wea78, Bow78, Wea79]. microfiche [Bei74]. microprocessor [Irb77]. Microprocessors [Lin78]. microprogrammed [Vis78]. mini [MH75, Sto78]. mini-language [MH75]. Minicomputer [Lew74, Sto77, MT73, TL79, WP77]. minimal [TL79]. minor [Lin79, Son75]. minority [Ano76, Ing75, Jon75a, Lew75, Lew76]. mirror [GG73]. MISS [SBS76]. mixed [MS76a]. mode [WA74]. Model [Fai78, Hed72, Mus76, Pow78a, Sal76]. Modeling [LK76]. Modelling [Pos79, BB76b]. models [KS76, MM78b, Sch73]. modified [PW75a]. modular [Ros76]. module [Swi76]. monitor [IN76]. monitors [Boy76]. Motivating [DS75]. motivational [Beu75]. motor [KPM74]. moves [Mou74]. MST [MB76a]. much [DP79, Sob76]. Multi [Rag76]. Multi-disciplinary [Rag76]. multifacetted [MK72]. multimedia [Han78]. multiprogramming [Bau76, WP77]. MUSIC [CJPT77]. my [Ham75c].

N [Jor72, Tho79]. nation [Any78]. National [BMOW76, Bat73, BKDH79]. nations [Rob75]. Native [Rot78]. natural [Cab73, Fre76a]. Naval [Wu74]. near [Har77]. Nebraska [NW78]. Need [Jon75a, CM75, Mou79, Oli73]. Needs [Fre76c, CAO^{+}76, God74, Rin78, Sch78b, Sta72, Wai79]. Network [Nie70a, BJ76, Hon73, Lew76]. Newtruck [Cla79]. Nirvana [Cur75]. non [Coo77b, Eck76b]. non-majors [Coo77b]. non-programming [Eck76b]. notations [Cou75]. note [Jal78, RD77]. Notes [Jär77, McF75, BD74a, BD74b, Bon75]. Notre [Sch71]. NSF [Lew76]. Numerical [Ort74, Cor78, Gin75].

O [Bai72]. objectives [Fre76c, Kei76, Bal78]. observations [Aik72, Tra76]. obtain [HFL^{+}78]. offering [Ell73]. offerings [Con75]. Ohio [Dal79]. Oklahoma [Hed77]. OLGA [TL79]. olin [KR76]. Olympia [Aik78b]. On-line [Che76b]. One [Har77, Bar75a, SW74b]. one-quarter [Bar75a]. Only [PS78b]. open [CMT74]. Operating [Cou77, Bau75, CKP78, CD79, Dal78c, Dun74, GG77, Gyl78, HP76, JK77, Lam76, Mac74, PW75a, RD77, Rus77, Tra74, WH73, WP77]. operational [Cab73]. operations [MPS78]. Opinions [Mou78]. opportunities [BP77]. option [WZ73]. ORACLE [HHP77]. oral [Cal71]. Oregon [Mou74]. organization [Bei73, Den70, Din74, MM73, PK74, Tha79]. organizations [Sch73]. oriented [BCS79, Che76c, Col79, Con77, DFF^{+}78, Gla75, Hon75, Jon75b, Mas73, Mic76, Mil75, PB74, Pow79, Sal77, Sen74, SI79, Tha79, Fai74b]. other [God73]. Ouchless [Bai72]. outline [Hof74]. outlines [Cav75]. overlooked [FK73]. overview [ES78, Mor76].

Pace [Var79a]. paced [Coo76, Lin76, Pal76]. package [BC70, Rag79, Tho79]. Packing [Fre76b]. paging [Lei77]. Panel [Ano76, BFC79, DAB^{+}79, EAAD79, FBB^{+}76, GKC^{+}74, Ham79, HLEA79, Maz79, Met79, Mou79, PDS^{+}79, ACEH76, Bon74, Chu75, McF75]. paper [LHL79, LAS^{+}77, Ric75, Sol78, Sto74b]. papers [Dal79]. paradigm [BB76b]. parallel [Kon75]. parameters [OWE73]. parametric [WH73]. part [DKMS75, HWC75]. partial [Har77]. parts [Bau76]. PASCAL [Bau79, TT78b, Lec74]. pays [San78]. PDP [TL79]. PDP-11 [TL79]. pedagogic [FK76]. pedagogical [Fle74, Hed72, Rei73, SR76]. pedagogy [Sen74]. perfect [Nor74]. performance [CGS75, Nie70b]. personal [Vaw76]. personalized [KW78]. perspective [NW78]. Ph [CC72, Mau71]. PH.D [Dep73, Wil72, Cal71, Tra76]. Ph.D. [Cla74, Cor70, Fin70, For70a]. phase [KPM74, MK74]. phase-locked-loops [KPM74]. philosophy [Ham75c, MK73]. pi [Bat73]. pilot [Dal77]. Pitt [Yel78]. PL [AIS72, AT77a, ES78, HW74, Lev77, NS76]. PL/1 [ES78]. PL/I [AIS72, AT77a, HW74, Lev77, NS76]. place [Bei77]. Plagiarism [Ott76]. Plan [MK74]. Planned [Luc76]. Planning [Bar75a, Jär77, MB76b]. plea [Lem78b]. PMS [Cou75]. point [Cab73]. pointer [Gum77]. pointers [CN76]. policy [Whi72]. policy-making [Whi72]. political [Ste72, Swi76]. Polytechnic [GG73]. poor [Pow78b]. POPSS [Cou77, WH73]. portability [NW78]. Position [Lov74, Ort74, Ric75, Sol78, Sto74b]. post [MS76a]. power [Eck76a]. Practical [TB77, Men74]. practice [Cha78a, FHSW78, Lev77, LK76, SW79]. practicing [FAT78]. Practicum [LT75, Web75, Ell75]. practitioners [Hof79]. pragmatic [FM78, FM79, WP78b]. pragmatically [Jon75b]. pre [Mou74]. pre-college [Mou74]. precollege [Mou78, Mou79, PDS^{+}79]. Predicting [CGS75]. predictions [New75]. predominantly [MS76b]. preferences [Maz76]. prejudices [Ham75c]. prelim [Mau71]. Preliminary [BKDH79, Ham79]. preparation [Arn78, Atc73]. preprocessed [Dal78b]. present [DMH^{+}77]. presented [McF75]. presenting [Tra74]. Prevention [Ott76]. Principles [Den70, Noo79, Sch78a]. privacy [Hof74]. probabilities [Vaw76]. Problem [Der70, MB77, Weg71, AC72, CM75, FK77, Pec77, PB77, Sen74, Woo78, HGU79]. problem-solving [CM75]. Problems [Aik71a, Ano70a, Ano70b, Ano71, For70b, Lav78, Lio70, Poi79, Rei72, Rog78, Roo72, BS76, Gra70, Rob75, TP72]. procedure [Lev77]. procedures [Den79b]. proceedings [Lit77]. Process [HGU79]. Processing [ACM72, Abb75a, BP77, BR79, Hag76, HT73, Kea75, KP76, KS78, Man74, MK73, Phi76, Rot78, Wes72]. processor [Bei73]. Produce [Bae75]. production [TB77, Tra76]. profession [Mil73]. professional [Aus74, BR79, Fre76a, JRS78, MM78a, PE73]. Professionalism [Aik71b]. professionals [Har74]. professors [Kor70]. Profile [Kal74, DAB^{+}79]. Program [AT77b, LSL^{+}76, Lov75, NW78, Nei76, Wel76, AP71, Aik71a, ABB^{+}77b, Bau70, Bei79, BCS79, BR79, Ber79, CC72, Che76a, Cla74, Col79, Dep73, FJ78, GM79, Ger75, Gil76, Hag76, Ham79, Hel73, Hom78, Hon75, Iye76, Iye78, Jon75b, JW78, KHKD78, Lak76, LL78, LI76, LAS^{+}77, MM78a, MK72, Mat70a, Mau77, MB76a, Mij78, Nar70, PB73, PB74, PC76, Rag79, RH73, Rhe73, RGH78, Sal77, Sap75, SCM78, Sem73, SW71, Tch75, Var79b, War75a, War75b, Web73, Wil72]. programme [CR70]. programmed [Boy76]. Programmer [Bar76, PW75a, DS72]. programmers [Cab73, FAT78, JW78, MB77, Sam71]. Programming [JW73, KP74, Ker78, Lev78, Oli74, PW75b, Sol78, Woo78, Ada75, Aik71a, Aik72, AHP77, AT76, Bai73, Bar75a, Bar78, Bas77, BT74, Bei73, Bei77, Bei74, Bez75, CN76, Cha74b, Cha78a, Che76c, CD79, Con74, Coo76, Coo77a, Coo78, CL76, Dal78c, DEN79a, DN75, DTW75, Der70, Dey73, DFF^{+}78, Din74, DMH^{+}77, Eck76b, Ell75, Emb78, ES78, FHW73, FM79, Fri75, FK76, FK77, FK79, Gra70, GKC^{+}74, Gri74, HSW78, Hon73, Hug77, IE72, Iye78, Jal78, Kei76, Kha76, Kha77a, KR72, Kim77, Kon75, Lan75, Law74, Lec74, Lem78a, Lem79, LAS^{+}77, Lov72, LM76, Mac79, Mac74, MF79, Mav76, McC78, Miz76b, New75, Noo79, Oml78, Pag78, PW74, PS78c, Rag76, Rag79, Rei73, Rip75, Sal73, Sch78a, Sch73, Seb76, SW74a, Shn77, SS74, Smi76]. programming [Som77, Tay77, Tro79, Wag75, WW74, Wei73, Wei78, Wil76a, vDSMM74, vDDDN76, AM79, LSL^{+}76]. Programs [Bae75, ACM72, Ano76, BMOW76, Cur75, DP79, God74, HLEA79, HAB^{+}76, JRS78, Kea75, Man74, Mei77, Pee76, Pow78a, Sal76, Sch79, Zel72]. progress [Boy76, DEN79a, LL78]. project [Cla79, Cla76, CKP78, Dal77, KS78, Lev78, Lin76, Mac74, PW75a, PS78a, PL78, Pow78b, Rus77, Ste72, Tha79, Vis78, Yel78, vdR79]. project-oriented [Tha79]. Projects [Roo72, Aik71a, Ano70a, Ano70b, Ano71, Cre78, For70b, KR79, Kru78, Lio70, Moc78, Rei72, Seb76]. promoting [Aik78a]. Proofs [Har78]. proper [WP78a]. proposal [BKDH79, Lin79, Map73]. proposals [Eck76a]. Proposed [Arc75, Sam71, CAO^{+}76, Hof79, JTF78, Pec75, Phi76]. prose [JW73]. prospective [God75]. prospects [deG72]. Provide [Hug70]. Providing [Wes72]. pseudo [Sjo76]. pseudo-assembler [Sjo76]. PSI [Mod74]. Public [KRFS79, BB79, God75, Man74]. publishers [Shn77]. Purdue [AM79]. purpose [Eck78, Lee72]. purposes [CJPT77]. pyramid [Eps75].

qualifying [Fin70, For70a, Nie70b]. quarter [Bar75a, LT74]. quasi [FL76, Bal78]. Quasi-Behavioral [Bal78]. quasi-interactive [FL76]. question [Wol76a]. questions [Fin70]. QUICK [FL76]. QUILT [Eck78].

rationale [ES78, FM78, FM79, Lee72]. Re [HA71, Sch77]. Re-evaluation [HA71]. RE-IPL [Sch77]. reaction [Bon74, DG74]. Reading [Kim79, LT76]. Real [Kru78]. Real-time [Kru78]. Realism [Fre76b]. realistic [Arn78, BS76]. rearview [GG73]. reason [Bre77]. reasons [Nei77]. recommendation [Kea75]. Recommendations [Dun74, HAB^{+}76, ABB^{+}77b, Ban78, Ham79, LSA^{+}76, LAS^{+}77, Rhe73, Wor78]. Recommended [Dav77, SLC^{+}74]. reconsidered [Che76c]. Record [Ham75a]. recruiting [BC79]. recursion [Ano77, AT76, BG74, Jac76b]. Recursive [Kim77, Kra72, Pee76]. redistricting [Ste72]. Reflections [Wax75]. regional [Mod73]. Reilly [Jor72]. related [Der70, JRS78, Lei72]. Relating [God74]. relationship [Aik70]. Relevance [CAO^{+}76, Aik71b, Shn71]. relevant [Miz76a]. Remarks [Sch71]. remedial [Din77, Tay79]. Remote [deG72]. Repackaging [EB79]. replaceable [Bau76]. reply [Kor75]. Report [Cai75, Hug70, Ing75, LSL^{+}76, Lit77, ABB^{+}77b, BFC79, DEN79a, Eng71b, Eng74, Han78, LR73, LL78, LSA^{+}76, MK74, Wak78, Wil78]. Representations [Bae75]. requirement [Cal71]. Requirements [DP79, Arc75, Miz76a, Zel72]. requisites [SI79]. research [Kay74]. resources [DKMS75, HWC75, War75b]. responsibilities [Des72]. responsibility [Kay74, Nie72]. Results [Kra72, FAT78]. returning [Dal77]. Review [Jor72, Kea75, SS79]. reviewing [Pow78a]. revision [Eng74, Eng76]. revolution [Hob78]. Riedinger [Jor72]. Rivest [BB79, SP79]. Rochester [BP77]. role [Aus74, Ell73, Fai74a, Har74, Hos73, Maz79, Pag78]. roots [Wei78]. RSA [Ber09]. rural [Shu79]. sabbaticals [Kor70]. SAM [Cav75]. SAMOS [GS75]. Sample [Mac79]. San [Haa78, Nei76]. Saudi [Air76]. scale [CR70, WW74]. schedule [Sha73]. Schenectady [Jor72]. school [Aik72, Atc73, BM75, BFC79, CS77, FK73, God75, KH75, MN76, Mus76, NT70, NFB^{+}78, Poi79, Tho75, Wol76b]. schools [Cha74a, Mou77, Shu79, KRFS79]. Science [Abb75a, BMOW76, CM75, CMT74, Dal79, God74, GH79, Haa78, HBJ76, Jor72, Kal74, LT75, MK74, Nei76, PB74, Sam75, ACM72, War75b, Abb75b, Abs78, AH72, AIS72, Agr76, Aik71b, AC72, Ano76, Any78, Arc75, AP70, Ash73, Atc73, AE72, Aus74, ABB^{+}77b, BC79, Bar75b, BP73, Bat73, BM75, BG76, Ber79, Beu75, Bon74, BNS78, BC70, BS76, BFC79, CK79, Car78, CC72, Cha74a, Cha77, Cha78b, Che76a, Chv78, Cla79, Cla74, Con77, CAO^{+}76, Coo77b, CR70, Cur75, Dal77, Dav77, DP79, Dep73, Der78, Des72, Din75b, DG74, EG76, Eck72, Ega76, EDL70, Eng70b, Eng70a, Eng71b, EB72, EJ73, Fai74a, FAT78, Fle74, FHSW78, FGH^{+}70, For70a, FJ78, FK73, FJ77, Gib77, GLO74, God73, Gol71, GS76, GG73, GB73, Gra78, Gro75, Ham75a, Ham79, HLEA79]. science [Ham70, HAB^{+}76, Ham75c, Han78, HM75, Har74, Hay74, HA71, Hel73, Hob78, Hom78, Hon75, Hos73, Hun70, Iye76, Iye78, Jac76a, Jeh74, JBF76, JRS78, Jon75b, KK78, Ker78, KL78, Kim78, KH74, KS74, LR72, LR73, Lav78, Lev79, Lev76, Lew76, LI76, Lin78, Lin79, Lov74, Lov72, Lus77, MF79, MM78a, MS77, Mal72, MK72, Mat78, Mat70b, Maz79, Mei79, MB76a, MB79, MM78b, Moc78, Mod73, Mod74, Moo70, Mou76, MN76, Mou78, Nah79, Nan76, New73, Nie70a, Nie70b, NR73, OR77, Oli73, Oml78, Pag78, PS78a, PE73, PB73, PB77, Poi76, Pol78, Pos79, Pow79, RH73, Ral71, Rin78, RGH78, Rob75, Ros76, Rot72, Rub76, Sal77, Sam74, San78, Sch78a, Sch79, Sch77, Sch71, Sem73, Sha72, Sha73, Shn71, Sjo76, Smi78, SR76, Smi79]. science [Smo76b, Sny70, Sol79, Som77, Son75, Son78, SG78, Sta72, SP70, SAK^{+}78, SW74b, Sti73, Sto74a, SW71, SW79, TP72, Tch75, TT78b, Tho74, Tho75, Tra76, TM74, Vic71a, Wal73, Wal74, War75a, Wea78, Wea79, WZ73, Weg70, Weg71, Wei77, Wil76b, WP78a, Win77, Wor78, YGM73, Yel78, Eng71a, Nei76, Pec75]. science/systems [MF79]. sciences [ABB^{+}77b, BL76, GS76, Lak76, Lei72, LAS^{+}77, Oli74]. scientist [Kay74]. scientists [Car78, Ham74, SI79]. scored [Boy76]. screen [Che76b]. search [Cor78]. second [HSW78, Ker78, MB79, Noo79, Pol72, Yel78]. Secondary [Mus76, Atc73, BFC79, NT70, PC76, Poi79, Wol76b]. security [Hof74]. Selecting [SR76]. selection [Sto77]. Self [Fre76a, Coo76, Lin76, NFB^{+}78]. Self-assessment [Fre76a]. self-explanatory [NFB^{+}78]. self-paced [Coo76, Lin76]. semester [ES78, Tho75]. seminar [AC72, Ell75, Nel71]. seminars [Fle74]. senior [Aik71a]. separate [EB79, GS76]. Separation [FHW73]. sequence [ASW78, ES78, HW74, Pow79]. Service [Sto74b, AH72, Bra70, God73, Jac76a, MM73, Mit79, Poi78, Rad76, Sto74a]. services [Lew75, Sto75]. session [ACEH76, Chu75, Ing75, Jeh74, Zin70]. Setting [Hon75, TB77]. sexual [Maz76]. Shamir [BB79, SP79]. shear [OWE73]. short [Ell73]. should [Gri74, HM75, HFL^{+}78]. show [Har78]. side [AP70]. SIGCSE [Dal79, Bon74, Hug70, Ing75, Lit77, McF75, Pol72, Wil78]. SIGCSE/CSA [Wil78]. Simon [SW74b]. simple [Bar78, BT74, Lev77, Pum75]. simulated [CJPT77]. Simulation [PL78, Aik72, JK77, KPM74]. Simulator [Gla75, AS78, Bei72, Lei77, MS76a, PK74, WH73]. simulators [SW74a, SP74]. single [Lev79]. ‘Sister [Aik78a]. six [MS76b]. SJCC [Zin70]. slides [MPS78]. Small [Gri78, AE72, CK79, Cha78a, DRT75, Din75a, Eng70b, Eng71b, FJ77, MHA78, Smi79, Weg71, Zim71, dC70]. Social [Nie72, Sam75, Shn71]. socialization [Swi76]. societies [Aus74]. society [ACEH76, Bat73, Dal76, HH73, Lee72]. Software [Bar75b, Bat74, Fai78, Fre76c, KR79, PB74, AHP77, BM78, Din77, Gra70, Hof78, Hof79, Irb77, JTF78, MB79, Par72, Sal77, Sob76, Son78, TB77, TMWD77]. solution [Har77]. solutions [Gra70, Poi79]. solve [Pec77]. Solving [HGU79, CM75, FK77, PB77]. Some [Eck76a, Fai74a, FK76, Hos73, Nei77, Noo79, Tra76, Web73, Weg70, Aik71a, Bra70, Cha75, TP72]. something [Har78]. sorting [MB76b]. soup [Mod74]. Space [Zel72]. special [Eck78, Ing75]. specialization [AP71]. specification [Bal78, Mei77]. Speech [SBS76]. speed [KPM74]. spread [Wes72]. Squanderland [Car78]. squares [Nor74]. staffing [DKMS75, HWC75]. stand [Bow78]. stand-alone [Bow78]. Standards [GCS77, Bre76]. Stanford [For70a]. Starter [Ran75]. State [Aik78b, GG73, Hed77, Lew76, LI76, Nei76, Nel71, HT73, Man74, Pec77, Fin70]. statement [Lov74]. Staten [CMT74]. States [Ham75b, Ham77]. Statistical [OWE73]. Statistics [Ham74, Wol76b]. Status [HLEA79]. stepwise [Seb76]. storage [AT77a, MHA78]. strategy [Kha77b]. Structure [God73, Mil73, Pal76, Pos79, Thu79, Var79b, Web73, Yel78]. Structured [Kon75, Lec74, Lev78, Miz76b, Oml78, Wag75, vDDDN76, AT76, Bar75a, BT74, Bei77, Bez75, Cha74b, CL76, Ell75, FK76, FK77, GKC^{+}74, HBJ76, HW74, Hug77, Jal78, Kei76, Kha76, Kha77a, Lev77, MH75, PB77, PW74, TB77, Wei78, Win77, WM74]. Structures [Iye76, AT77b, Ber76, DRT75, Din74, EJ73, Eps75, MF79, Pra76, Ric75, SB74, TM74, TS75, Wax75, YGM73, BM78]. Student [Bai73, MM73, Wil73, Arn78, Bau70, Boy76, BB76b, CKP78, DAB^{+}79, DP79, GH78, Har77, Hon73, Lin76, LK76, Mas73, Mil75, RD77, SCM78, Sch77, Smo76b, Ste72]. student-computer [Hon73]. student-designed [Bau70]. student-operating [RD77]. student-oriented [Mas73]. student-use [GH78]. Students [Rot75, Aik72, Bre76, CGS75, Dal77, DEN79a, DN75, DS75, Gol71, Mat78, NW78, PC76, Rot78, SW74a, Shn77, Smi78, Son75, Ste78, Wei77, Wes72]. studies [DAC^{+}78, Lei72]. studies-implementing [DAC^{+}78]. study [Agr76, Aik71b, GG78, KR72, Mor72, RD77, Wol76a]. studying [HP76]. Style [KP74, Fre76b, Smo76a]. Sub [LSA^{+}76, LAS^{+}77]. Sub-Committee [LSA^{+}76]. Subcommittee [LHL79]. subjectively [Nah79]. subsets [HW74]. substitute [BP73]. subsystems [Thu79]. successful [SCM78]. suggested [Sap75]. suggestion [Sch77]. supply [Cab73, Ham75b, Ham77]. support [Ano76]. supported [Jon75a, Mor72]. survey [Abb75b, Che75, FAT78, FJ77, Hof78, Hom78, Lem79, Rei73, Rub76, Sch79, Wai79, Wal73, vDSMM74]. Syllabi [For70a]. Syllabus [Nie70b, Cor70]. symbol [Fin70]. Symbolic [CC72]. Symposium [Dal79, Ing75, McF75, Pol72, Wil78]. syndrome [Har77]. syntax [BG74, Hed72]. synthesis [SBS76]. system [Bau75, Bau76, Che76b, Chu75, CKP78, Dal78c, DTW75, Den70, DFF^{+}78, FL76, Fre76a, GH78, Gil76, Gri78, GG77, Hon73, KP76, KF76, Lam78, LT74, LT78, Mar72, Mic76, Pal76, RD77, SBS76, SB76, Tra74, Vic71b, Vis78, Wak78, WH73, WW74, WA74, WP77]. Systematic [Bar78]. Systems [Bae75, EDL70, SG78, Ash73, BCS79, BS76, Cai75, Cha78a, Cla74, Cla76, CD79, CL76, Cou77, Dun74, FBB^{+}76, Gra70, Gri78, Gyl78, HP76, Jär77, JK77, Lam76, Lan75, Mac74, MF79, Map73, MT73, Mau71, Mij78, Mor76, Pec77, PW75a, Roo72, Rot75, Rus77, Sam74, SHLS79, TMWD77, Vis77, WW74, Win78]. Számok [Bar76].

Tab [Wel76]. tables [JW73]. tailored [Whi77]. tasks [Nei77]. teach [Bei74, Gri74, Lam76, MPS78, Vis77, Vis78, Wea79]. Teacher [HJ73, Moc78, BM75, Poi76]. Teacher-training [Moc78]. teachers [Abs78, Atc73, BFC79, FK73, Jac76a, Mal72, Mar76, McG73, MN76, Mou78, Mou79, NT70, Shu79]. Teaching [Ada75, App79, Bez75, CMT74, EW72, Emb78, FK77, Gra70, Gyl78, Irb77, Jon79, Lem79, LT78, Lov79, LM76, Mar76, Mat74, McC78, NS76, PW74, PK74, Sal77, SP74, Tay77, Var79b, Wol76b, Aik72, BM78, BC77, CJPT77, Car78, CN76, CM75, Cha74b, CL76, Dal78d, DEN79a, Dey73, Eps75, Fri75, FK76, Ger75, GS76, Hon73, Hun70, IE72, Jal78, Kha77b, KF76, Lan75, Lec74, Lem78a, Lev77, Luc76, MF79, Mau77, Mav76, MD76, MH75, MK73, New73, NW74, OR77, Rad76, Rog78, RJB71, San78, SM76, Sha73, TT78b, Wu74, YGM73]. teaching-machine [NW74]. Team [Cre78, HHD78, PW75a, KR79, KS78, Lem78a]. Technical [Dal79, HT73, McF75, Pol72, Wil78]. technique [Jac76b]. Techniques [Abs78, TT78a, Bei72, Cla76, Fle74, LM76, New73, Par72]. Technological [Whi72, Sam71]. technology [Aik78a, BL76, Maz79, BP77, MK74]. television [Bai73, Dal76]. ten [Sch78a]. tenth [Dal79]. term [Cla79]. terminal [Gla75, PW75b]. test [Boy76, Mus76, Pal76]. testing [Tro79]. text [Wol76a]. textbook [Nel78]. textbooks [Shn77]. their [FAT78]. them [Shu79]. theoretical [Gra78]. theory [Aik70, FHSW78, Har78, HSW78, Hon73, Swi76, SW79, Vaw76]. Thermodynamics [Wu74]. thought [Kay74]. thoughts [Aik71a, Fai74a, Nei77, Web73, Weg70]. Three [Mas73, Lit77]. TI980A [Bau75]. TICCIT [Mor76]. time [Kru78]. tool [BM78, HHP77, HP76, Hug77, Lev77, MB76b, TT78b]. tools [FM79, Pol78, SR76]. top [GG77, Hed72, Pum75, Rag79, TB77]. top-down [GG77, Hed72, Pum75, Rag79, TB77]. topics [Dav77]. Topological [MB76b]. Toronto [Cor70]. Trace [BBS75]. track [SW74b]. traditional [FJ78]. Training [Mor77, Dal77, HJ73, Lak76, McG73, Moc78, Pec77, Rot75, Tay79, Wil76b]. transfer [Aik78a, Maz79, Mil75]. transformations [MD76]. translator [Kra72]. translators [Cal78]. transparencies [Bei74]. traveling [Nel71]. traveral [Den79b]. tree [Den79b]. trends [JRS78]. trigonometry [SS79]. Turning [Sch77, MS76a]. turning/post/mixed [MS76a]. tutor [DN75, NW74, Lin76]. tutorial [SS79, Wu74]. tutorials [Lit77]. tutoring [BB76b]. TV [EG76]. Two [Bae75, DAC^{+}78, AM79, BG76, Cab73, Con75, ES78, HT73, Kea75, MM78b, Pec75, PB73, Sti73]. two- [HT73]. two-semester [ES78]. two-year [Cab73, Con75, Kea75, Pec75]. types [Gum77].

U.S. [Slo74, Wu74]. U302 [Roo72]. Undergraduate [Ham70, Hay74, TP72, ABB^{+}77b, Bau70, Cha75, Che76a, Cla76, CR70, Cre78, Dal78a, DP79, Der78, Dow74, Eng70a, EJ73, FS79, Gin75, Gri78, HAB^{+}76, HHD78, HBJ76, Hom78, Lam78, Lov72, Lus77, Man74, MK72, PS78a, Rhe73, Rus77, Sam74, Sap75, Sch77, Som77, SB74]. undevelopment [Tch75]. unfort [Emb78]. unifies [FHSW78]. unique [LI76]. United [Ham75b, Ham77]. Universal [SP79, SP70, SW74b]. Universities [Hug70, Chv78, JBF76, Lew76, Man74]. University [AM79, Eck76a, Fin70, Hay74, Lew76, LI76, MS76b, Noe73, AP70, BG76, CGS75, Ega76, Mod73, Pec77, SB76, Var78, Cor70, For70a, GG73, Haa78, Hag76, Hed77, Hun70, NW78, Nei76, Nel71, Var79a, Web75, WP78b]. up-to-date [FAT78]. update [CAO^{+}76, Ham77]. updating [Sam71]. upgrade [Oli73]. upper [BG76]. upside [Wil72]. Upsilon [Bat73]. usage [Oli74, PC76, Wai79]. Use [For73, HBJ76, Pow78a, Bau75, Bau76, Bei74, CN76, GH78, Gla75, Han78, Kra72, Maz78, McG73, Pal76, Pol78, Smi76, TT78b, Tra74, Var78, Vaw76, WM74]. user [BJ76, Che76c]. user-oriented [Che76c]. uses [KS76, KR76, Mor77]. Using [Lam76, Lam78, Lev79, MPS78, Wea79, Wor78, Bar75a, Coo77a, EW72, ES78, FK76, Irb77, Kei76, LT78, Lov79, MHA78, RD77, WW74]. utility [JK77]. utilization [Coo78, DKMS75, HWC75]. utilize [Shu79]. utilizing [Mil78]. UW [GS76].

V [HP76]. VACS [Gro75]. valuable [Mod74]. variables [Gum77]. variety [Mor76]. vehicle [PL78]. verification [Ger75, JW78]. versus [Lem78b]. via [Emb78]. Video [Gro75, Mic76]. view [Con73, Mar75]. Virginia [GG73]. Visiting [Kor70]. Visual [KHKD78, SW74a]. voie [San78]. vs [Cal78, Pec77].

want [Mou79]. Washington [Aik78b, Hun70]. Waxman [Kor75]. way [DMH^{+}77]. Wegner [DS72]. Western [Moc78]. Wheaton [LR73, LR73]. which [Bae75, Bei72]. while [Luc76]. Whither [Hos77]. who [Shu79]. wide [Wes72]. Wisconsin [WP78b, BFC79]. without [Ano77, DEN79a]. Women [Dal77]. work [HA78]. workable [Lev78]. working [ABB^{+}77b, LHL79, LAS^{+}77]. workshop [LR73]. writing [Bar75a, Bas77, Dal78d, Din75a, HHP77, Nel78]. written [Cor70].

XS [NFB^{+}78]. XS-0 [NFB^{+}78].

year [AM79, BG76, Cab73, Con75, DAC^{+}78, HT73, JTF78, Kea75, Mas73, Pec75, PB73, Smo76b, Sti73]. York [Fin70]. Young [Hay74]. yourself [Abb77a].

Zealand [Moo70].

References

Abbey:1975:DPC

[Abb75a] Duane C. Abbey. Data processing and computer science graduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):71–75, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Abbott:1975:ISC

[Abb75b] Russell J. Abbott. An informal survey of computer science coursesr. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):84–86, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Abbott:1977:DIY

[Abb77a] R. J. Abbott. A do-it-yourself instant compiler.kit. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):53–58, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Austing:1977:CRU

[ABB^{+}77b] Richard H. Austing, Bruce H. Barnes, Delle. T. Bonnette, Gerald L. Engel, and Gordon Stokes. Curriculum recommendations for the undergraduate program in computer science: a working report of the ACM committee on curriculum in computer sciences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(2):1–16, June 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on computer science curricula.

Abshire:1978:TCS

[Abs78] Gary M. Abshire. Techniques for computer science teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):42–46, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alter:1972:ECS

[AC72] Ronald Alter and Thaddeus B. Curtz. An experimental computer science problem seminar. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):144–153, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Austing:1976:CSC

[ACEH76] Richard Austing, William Cotterman, Gerald Engel, and Ellis Horowitz. Computers and society courses and computer literacy (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):324, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Staff:1972:CDP

[ACM72] ACM SIGCSE Bulletin Staff. Correction to “degree programs in computer science, processing, etc. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(4):45–46, December 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:1975:TDP

[Ada75] J. M. Adams. Teaching declarative programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):83–85, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Austing:1972:CSE

[AE72] Richard Austing and Gerald Engel. Computer science education in small colleges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):38, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Agresti:1976:CSI

[Agr76] William W. Agresti. Computer science as an interdisciplinary study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):12–14, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Adams:1972:ISC

[AH72] J. M. Adams and D. H. Haden. Introductory service courses in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):49–52, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alexander:1978:CEW

[AH78] Roberta R. Alexander and Frank J. Hatfield. Cooperative education: what is involved? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):96, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Alford:1977:SEA

[AHP77] M. Alford, P. Hsia, and F. Petry. A software engineering approach to introductory programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):157– 161, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Aiken:1970:CRF

[Aik70] Robert M. Aiken. A course on the relationship of formal language theory to automata. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):34–40, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aiken:1971:PEP

[Aik71a] Robert M. Aiken. Problems, exams, projects: some thoughts on the examples of program assignments for a senior level programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(3):19–24, September 1971. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aiken:1971:PRG

[Aik71b] Robert M. Aiken. Professionalism, relevance and graduate study in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(2):25–27, June 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aiken:1972:EOT

[Aik72] R. M. Aiken. Experiences and observations on teaching computer programming and simulation concepts to high school students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):67–71, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aiken:1978:SCA

[Aik78a] Robert M. Aiken. The ‘Sister City Approach’ — promoting transfer of technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(2):10–11, June 1978. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aikin:1978:CLI

[Aik78b] John O. Aikin. Computer literacy: an interdisciplinary, hands on approach at the Evergreen State College, Olympia Washington. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):8–12, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Aird:1976:CMD

[Air76] Charles L. Aird. Computer manpower development in Saudi Arabia. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):125–128, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Adams:1972:PCS

[AIS72] J. Mack Adams, William H. Inmon, and Jim Shirley. PL/I in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):116– 126, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:1979:TYC

[AM79] A. J. Adams and John Maniotes. The two year curriculum in Computer Programming at Purdue University Calumet campus. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):207–213, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Andrews:1972:CY

[And72] W. B. Andrews. The computer and you. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):103– 109, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:1970:PEPa

[Ano70a] Anonymous. Problems, exams, projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(1):15– 18, March 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:1970:PEPb

[Ano70b] Anonymous. Problems, exams, projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(4):55, September/October 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:1971:PEP

[Ano71] Anonymous. Problems, exams, projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(1):40– 46, March 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:1976:FPS

[Ano76] Anonymous. Facilities and programs in support of education in computer science at minority institutions (panel discussion). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):94, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Anonymous:1977:DFD

[Ano77] Anonymous. Depth-first digraph algorithms without recursion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):151–153, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Anyanwu:1978:CSE

[Any78] J. A. Anyanwu. Computer science education in a developing nation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):37–40, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Armstrong:1970:BUC

[AP70] Russell M. Armstrong and Emmett K. Platt. Business and the university computer science department: the left-hand side of a dialogue. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):6–8, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Abrams:1971:MPS

[AP71] M. D. Abrams and J. H. Pugsley. M.S. program with specialization in computers: department of electrical engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(3):25–37, September 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Appelbe:1979:TCD

[App79] Bill Appelbe. Teaching compiler development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):23–27, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Archibald:1975:PMR

[Arc75] J. A. Archibald, Jr. Proposed mathematics requirements for the Bachelor’s degree in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(4):12– 15, December 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Arnow:1978:CER

[Arn78] Barry Arnow. Cooperative education and realistic student preparation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):86, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Albright:1978:BML

[AS78] Richard A. Albright and Bruce D. Smith. BRAINO: a machine language simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):59–65, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ashenhurst:1973:ICS

[Ash73] Robert L. Ashenhurst. Implications for computer science departments of the ACM information systems curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):2–5, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Ahlgren:1978:SCC

[ASW78] D. Ahlgren, A. Sapega, and H. Warner. A sequence of computing courses for liberal arts colleges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):180–182, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Augenstein:1976:LRS

[AT76] Moshe Augenstein and Aaron Tenenbaum. A lesson in recursion and structured programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):17–23, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Augenstein:1977:ABS

[AT77a] Moshe Augenstein and Aaron Tenenbaum. Approaches to based storage in PL/I. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):145–150, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Augenstein:1977:PED

[AT77b] Moshe Augenstein and Aaron Tenenbaum. Program efficiency and data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):21–27, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Atchison:1973:CSP

[Atc73] William F. Atchison. Computer science preparation for secondary school teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):45–47, February 1973. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Austing:1974:RPS

[Aus74] Richard H. Austing. The role of professional societies in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(4):15–20, December 1974. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baecker:1975:TSW

[Bae75] Ron M. Baecker. Two systems which produce animated representations of the execution of computer programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):158–167, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Graphics/imager/1975.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Graphics/siggraph/

1975.bib. Proceedings of the 5th SIGCSE symposium on Computer science education.

Bailey:1972:O

[Bai72] T. E. Bailey. Ouchless I/O. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(3):8–18, October 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bailey:1973:SAT

[Bai73] T. E. Bailey. Student attitudes toward a new application of instructional television in a computer programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):173–176, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Baldwin:1978:QBO

[Bal78] Lynne J. Baldwin. Quasi-Behavioral Objectives for curriculum specification. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):1–7, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Banes:1978:IAC

[Ban78] Bruce Banes. Implementation of the ACM curriculum recommendations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):177, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Barnard:1975:PEO

[Bar75a] A. C. L. Barnard. Planning and experience with a one-quarter course on compiler writing using Gries’ book and structured programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):27–29, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnard:1975:SEA

[Bar75b] A. C. L. Barnard. Software engineering in an academic computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):62–70, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bardos:1976:PME

[Bar76] A. F. Bardos. Programmers’ mass education at Számok. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(2):41–44, June 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnett:1978:SIS

[Bar78] Michael P. Barnett. Systematic instruction in simple programming gambits. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):108–112, August 1978. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Bash:1977:WAP

[Bas77] Joyce M. Bash. The writing of an assembler in the programming language course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):88–90, December 1977. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bateman:1973:UPE

[Bat73] Barry L. Bateman. Upsilon pi epsilon: the national computer science honor society. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(2):23–25, June 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bateman:1974:SC

[Bat74] Barry L. Bateman. Software and Curriculum 68. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6 (3):88–89, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bauer:1970:SDU

[Bau70] Michael A. Bauer. A student-designed undergraduate program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):13–17, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bauer:1975:DTO

[Bau75] Henry R. Bauer III. The design of a TI980A operating system for classroom use. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):20–22, February 1975. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Bauer:1976:CUR

[Bau76] Henry R. Bauer III. Classroom use of replaceable parts in a multiprogramming system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):108–110, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Bauer:1979:EPI

[Bau79] Michael A. Bauer. Experiences with PASCAL in an introductory course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):158–161, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Barr:1976:IIB

[BB76a] Avron Barr and Marian Beard. An instructional interpreter for basic. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):325–334, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Burton:1976:TSM

[BB76b] Richard R. Burton and John Seely Brown. A tutoring and student modelling paradigm for gaming environments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):236–246, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE– SIGCUE joint symposium on Computer science education.

Blakley:1979:RSA

[BB79] G. R. Blakley and I. Borosh. Rivest–Shamir–Adleman public key cryptosystems do not always conceal messages. Computers and Mathematics with Applications, 5(3):169–178, 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Biermann:1975:TIA

[BBS75] A. W. Biermann, R. I. Baum, and M. Silverman. Trace information as an aid to debugging. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):44–49, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brillinger:1970:CPI

[BC70] P. C. Brillinger and D. D. Cowan. A complete package for introducing computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):118–126, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bremmer:1977:ETC

[BC77] Dale Bremmer and Bart Childs. An experience in teaching “Cobol?” to graduate engineers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):96–99, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Bailes:1979:CSE

[BC79] Gordon L. Bailes and Terry A. Countermine. Computer science (1978) enrollment, faculty, and recruiting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(2):43– 51, June 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Benjamin:1979:COG

[BCS79] James E. Benjamin, Wilbur L. Campbell, Jr., and John F. Schrage. Career oriented graduate program in management systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):10–13, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Bonnette:1974:ENa

[BD74a] Della Bonnette and John F. Dalphin. Editorial notes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(2):1, June 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bonnette:1974:ENb

[BD74b] Della T. Bonnette and John F. Dalphin. Editorial notes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(4):1, December 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beidler:1972:SWE

[Bei72] John A. Beidler. A simulator which emphasizes addressing techniques. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(4):11–18, December 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beidler:1973:MIC

[Bei73] John A. Beidler. A machine independent course in processor organization and assembler language programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5 (1):149–152, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Beil:1974:UTM

[Bei74] Donald H. Beil. The use of transparencies and microfiche to teach computer programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):51–59, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beidler:1977:PAS

[Bei77] John Beidler. A place for assembler in structured programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):108–112, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Beil:1979:CPM

[Bei79] Donald H. Beil. A course in program maintenance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(2):19–22, June 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berztiss:1976:WHD

[Ber76] A. T. Berztiss. The why and how of discrete structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):22–25, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Berztiss:1979:MPC

[Ber79] Alfs T. Berztiss. The M.S. program in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):61–69, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Bergmann:2009:DKR

[Ber09] Seth D. Bergmann. Degenerate keys for RSA encryption. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):95–98, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beus:1975:MAB

[Beu75] H. Lynn Beus. A motivational approach to basic computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):33–38, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bezanson:1975:TSP

[Bez75] William R. Bezanson. Teaching structured programming in FOR-TRAN with IFTRAN. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):196–199, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Burd:1979:RCS

[BFC79] Barry Burd, Timothy Fossum, and Donald Chambers. A report on computer science certification for secondary school teachers in Wisconsin (panel discussion). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):245–246, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Bell:1974:LRS

[BG74] Stoughton Bell and Edgar J. Gilbert. Learning recursion with syntax diagrams. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):44–45, September 1974. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berk:1976:CSC

[BG76] Toby S. Berk and Malcolm H. Gotterer. A computer science curriculum for a two year upper division university. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8 (3):15–21, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Beerman:1976:BEC

[BJ76] Thomas H. Beerman and Gregory S. Jonesku. BECUN: the educational computer user’s network at Battelle. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):182–188, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE– SIGCUE joint symposium on Computer science education.

Bork:1979:PPN

[BKDH79] Alfred Bork, David Kniefel, Karen Duncan, and Donald Holznagel. Preliminary proposal for a National Educational Computing Consortium. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(3):9–10, September 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bordage:1976:ICA

[BL76] G. Bordage and Robin B. Lake. An introductory course in the applications of computer technology in the health sciences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):184–188, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Bauer:1975:CSH

[BM75] Charles R. Bauer and John C. Meinke. Computer science for the high school teacher. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):19–20, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beidler:1978:STT

[BM78] John Beidler and John Meinke. A software tool for teaching Data Structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):120–125, August 1978. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Barnes:1976:NSF

[BMOW76] Bruce H. Barnes, Andrew R. Molnar, Lawrence H. Oliver, and Robert F. Watson. National science foundation programs in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):1, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Brackett:1978:CSE

[BNS78] Gail B. Brackett, Chadwick H. Nestman, and Emil R. Spees. Computer science education and the ideal of liberal arts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):13–15, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Bolz:1973:ICE

[Bol73] Harold A. Bolz. Impact of computers on engineering education (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):1, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Bonnette:1974:SPD

[Bon74] Della Bonnette. SIGCSE panel discussion, the industry’s reaction to computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(4):30–35, December 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bonnette:1975:EN

[Bon75] Della T. Bonnette. Editorial notes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):1, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bowles:1978:CCB

[Bow78] Kenneth L. Bowles. A CS1 course based on stand-alone microcomputers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):125–127, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Boyle:1976:CSP

[Boy76] Thomas A. Boyle. Computer-scored, programmed test monitors student progress. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):96–99, February 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Bateman:1973:CSF

[BP73] Barry L. Bateman and Gerald N. Pitts. Computer science as a foreign language substitute. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):132–133, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Bell:1977:EOD

[BP77] Donald H. Bell and Joseph W. Panko. Educational opportunities for the deaf in data processing at Rochester Institute of Technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):79–84, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Benson:1979:CBP

[BR79] Robert J. Benson and Robert A. Rouse. A community based professional development program in data processing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):136–139, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Brady:1970:ISC

[Bra70] Allen H. Brady. The introductory and service courses in computing: some experiences and a critical assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2 (2):31–36, June 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brewer:1976:DSB

[Bre76] Richard K. Brewer. Documentation standards for beginning students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):69–73, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Brewer:1977:MRK

[Bre77] Richard K. Brewer. Macrognosographics: a reason to keep flowcharting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):154–156, February 1977. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Buck:1976:IIS

[BS76] John Buck and Ben Shneiderman. An internship in information systems: Combining computer science education with realistic problems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):80–83, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Basili:1974:ESS

[BT74] Victor R. Basili and Albert J. Turner. Experiences with a simple structured programming language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):144– 147, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Bunt:1971:DA

[Bun71] R. B. Bunt. Dissertation abstracts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(4):46–52, December 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cable:1973:CET

[Cab73] Frank Cable. Computer education in two-year institutions: — A natural point of supply for hireable, operational, EDP-applications programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):32–37, February 1973. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Cain:1975:RDS

[Cai75] James T. Cain. Report of the digital systems education committee. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):13–16, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Calingaert:1971:OFL

[Cal71] Peter Calingaert. An oral foreign language requirement for the Ph.D. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(4):21–23, December 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Calingaert:1978:CCV

[Cal78] Peter Calingaert. Compiler course vs. translators course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):20–23, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Conti:1976:RCS

[CAO^{+}76] Dennis Conti, Russell Armstrong, Paul Oliver, Robert Orchard, and John Shoosmith. Relevance of computer science education to industry and government needs — A critique of the proposed update to curriculum ’68. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):92, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Carroll:1978:CCS

[Car78] John M. Carroll. The case for computer scientists teaching computer science or digit-Alice in Squanderland. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):229–231, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Caviness:1975:SCO

[Cav75] B. F. Caviness. SAM course outlines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):92– 103, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Caviness:1972:SMC

[CC72] B. F. Caviness and G. E. Collins. Symbolic mathematical computation in a ph. D. computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4 (1):19–23, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Conner:1979:AIG

[CD79] William M. Conner and Kenneth A. De Jong. The academic/ industry gap in systems programming and operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):118–126, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Capstick:1975:PPU

[CGS75] C. K. Capstick, J. D. Gordon, and A. Salvadori. Predicting performance by university students in introductory computing courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):21–29, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Coulter:1978:CC

[CH78] Neal S. Coulter and Fred G. Harold. Computer currency. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):156–161, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Chand:1974:CSE

[Cha74a] Donald R. Chand. Computer science education in business schools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):91–97, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chand:1974:SDL

[Cha74b] Donald R. Chand. A structured design language for teaching programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):121–124, February 1974. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Chanon:1975:CCU

[Cha75] R. N. Chanon. Compiler construction in an undergraduate course: some difficulties. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):30–32, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chanon:1977:EIC

[Cha77] R. N. Chanon. An experiment with an introductory course in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):39–42, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Chanon:1978:CPP

[Cha78a] R. N. Chanon. A course in programming and practice: toward small systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):224–228, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Chanon:1978:GIC

[Cha78b] R. N. Chanon. A graduate immigration course in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):32, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/ CSA Technical Symposium on Computer Science Education.

Chanon:1978:EA

[Cha78c] Robert N. Chanon. Educational AIDS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(2):66– 69, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cheney:1975:ICS

[Che75] Robert S. Cheney. An introductory computer survey course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):30–34, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cheng:1976:DUC

[Che76a] Richard Cheng. A diversified undergraduate computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):366–370, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Cheng:1976:LLS

[Che76b] Richard Cheng. On-line large screen display system for computer instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):179–181, February 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Cherniak:1976:IPR

[Che76c] Bob Cherniak. Introductory programming reconsidered — a useroriented approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):65–68, February 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Chu:1975:PSD

[Chu75] Yaohan Chu. A panel session on “digital system education”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):212, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Chvalovsky:1978:CSE

[Chv78] Vaclav Chvalovsky. Computer science education at universities: the case of developing countries. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):41–47, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Carpenter:1977:MSC

[CJPT77] B. E. Carpenter, P. C. Jenkins, L. W. Pearson, and L. K. Thomas. MUSIC: a simulated computer for teaching purposes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):70–76, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cameron:1979:CSC

[CK79] J. S. Cameron and Z. A. Karian. Computer science curriculua for small colleges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):215–219, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Comaa:1978:SGP

[CKP78] H. Comaa, J. Kramer, and B. K. Penney. A student group project in operating system implementation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):197– 202, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Cooper:1976:IHA

[CL76] Roger T. Cooper and Malcolm G. Lane. An improved handson approach to teaching systems programming and the impact of structured programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):115–124, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Clark:1974:HSC

[Cla74] Douglas Clark. Hardware systems in the core curriculum of a computer science Ph.D. program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):106– 110, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Claybrook:1976:CPD

[Cla76] Billy G. Claybrook. A course and project description for an undergraduate course in evaluation and measurement techniques for computer systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):111–114, July 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Clair:1979:NCL

[Cla79] Daniel C. Clair. Newtruck: a comprehensive long-term project in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):131–135, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Cashman:1975:NTP

[CM75] W. F. Cashman and W. J. Mein. On the need for teaching problemsolving in a Computer Science Curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):40–46, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Chi:1974:CSS

[CMT74] Emile C. Chi, Marsha Moroh, and Miriam R. Tausner. Computer science at staten island community college: Teaching computer science in an open admissions environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):48– 52, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Carvalho:1976:UPT

[CN76] Sergio E. R. Carvalho and Miguel Angelo A. Nóvoa. On the use of pointers and the teaching of disciplined programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):26–29, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Collins:1979:DOI

[Col79] William J. Collins. A data-oriented introduction to program design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(4):49–55, December 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Connolly:1973:CJC

[Con73] Frank W. Connolly. A community/junior college view of Curriculum ’68. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):68–69, February 1973. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Conway:1974:IIP

[Con74] Richard W. Conway. Introductory instruction in programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):6–10, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Connolly:1975:CTY

[Con75] Frank W. Connolly. Computers at two-year colleges: current course offerings and facilities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):6–11, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Conner:1977:COA

[Con77] William M. Conner. A concept-oriented approach to introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):162–164, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Cook:1976:SPI

[Coo76] Curtis R. Cook. A self-paced introductory Fortran programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):78–79, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Cook:1977:APC

[Coo77a] Curtis R. Cook. Applications programming course using guided design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):79–82, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Cook:1977:AIC

[Coo77b] Robert N. Cook. An approach to the introductory computer science course for non-majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):30–33, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Cook:1978:MUB

[Coo78] Robert N. Cook. Microcomputer utilization in basic programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):23–26, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Corneil:1970:SUT

[Cor70] Derek G. Corneil. The syllabus for the 1970 University of Toronto Ph.D. written comprehensive examination. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(2):41–43, June 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Corliss:1978:SIC

[Cor78] George Corliss. A search for interactive computer graphics lessons in numerical analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(2):17–20, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Coulter:1975:IPI

[Cou75] Neal S. Coulter. Introducing PMS and ISP notations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(4):72–75, December 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Coulter:1977:OSP

[Cou77] Neal S. Coulter. Operating systems with POPSS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):85–87, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Cowan:1970:LSU

[CR70] D. D. Cowan and R. B. Roden. A large-scale undergraduate programme in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):18–23, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Crenshaw:1978:TPU

[Cre78] John H. Crenshaw. Team projects in the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):203–205, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Costello:1977:GBS

[CS77] Donald F. Costello and Richard J. Schonberger. On guiding the business school toward computer literacy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):180– 183, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Curtis:1975:CSF

[Cur75] Kent K. Curtis. Computer science, federal programs, and Nirvana. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):109–113, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Dale:1979:CPD

[DAB^{+}79] Nell Dale, Richard Austing, Lynne J. Baldwin, Lucia McKay, and Mary Shaw. Curriculum 78 (panel discussion): Consequences for the student profile. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):87–88, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Dillman:1978:TYC

[DAC^{+}78] Richard W. Dillman, William H. Anderson, Dorothea L. Choper, John M. Lloyd, Katherine B. Simms, and James F. Williams. Two year curricula in computer studies-implementing the guidelines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):140–150, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Dalphin:1976:CSC

[Dal76] John F. Dalphin. A computers and society course: Adapted for television. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):98–100, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Dale:1977:WSP

[Dal77] N. Dale. Women in science: a pilot project in accelerated computer training for returning women students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):10– 15, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Dale:1978:DMU

[Dal78a] Nell Dale. Database management in the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):136–139, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Daley:1978:BPL

[Dal78b] James Daley. Basic2 — a preprocessed language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10 (2):70–73, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Daley:1978:ICP

[Dal78c] James Daley. An inexpensive concurrent programming language for operating system construction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):52–55, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Daley:1978:LAT

[Dal78d] James S. Daley. A laboratory approach to teaching compiler writing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):19–21, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Dalphin:1979:PTS

[Dal79] John F. Dalphin, editor. The papers of the tenth SIGCSE Technical Symposium on Computer Science Education, Dayton, Ohio, February 22–23, 1979, volume 11(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA76.27 .A79a vol. 11, no. 1.

Davis:1977:RMT

[Dav77] Ronald L. Davis. “recommended mathematical topics for computer science majors”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):51–55, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

deCampo:1970:ICS

[dC70] Leila de Campo. Introducing the computer at a small liberal arts college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):113–117, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

deGrasse:1972:RCH

[deG72] Richard V. deGrasse. Remote computing in higher education: prospects for the future. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):39–41, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denning:1970:PCS

[Den70] Peter J. Denning. Principles of computer system organization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):45–55, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Daly:1979:PRT

[DEN79a] Cecilia Daly, David W. Embley, and George Nagy. A progress report on teaching programming to business students without lectures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):247–250, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Denenberg:1979:ICB

[Den79b] Stewart Denenberg. Increasing the clarity of binary tree traveral procedures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(2):36–39, June 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dept:1973:CSP

[Dep73] CORPORATE Carnegie-Mellon Univ. C. S. Dept. The computer science PH.D. program at CMU. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(2):33–40, June 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dershem:1970:PRC

[Der70] Herbert L. Dershem. Problem related to card games for an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(5):48–49, December 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dershem:1978:ECU

[Der78] Herbert L. Dershem. An experience component for undergraduate computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):24–26, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

DesJardins:1972:RCS

[Des72] R. B. DesJardins. The responsibilities of computer science in continuing education for managers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):57–59, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dey:1973:IMT

[Dey73] Suhrit K. Dey. Inductive methods of teaching programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):6–8, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

DiManzo:1978:CCS

[DFF^{+}78] M. Di Manzo, M. Ferraris, A. L. Frisiani, V. Midoro, and G. Olimpo. COMES: a CAI system oriented to programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):7–15, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Dodd:1974:IRC

[DG74] George Dodd and Fred A. Gluckson. Industry reaction to computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):79–80, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Dinerstein:1974:ECP

[Din74] N. T. Dinerstein. An experimental course in programming languages, data structures and machine organization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):30–32, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dinerstein:1975:CWC

[Din75a] N. T. Dinerstein. A compiler writing course in a small college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):71–72, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dinerstein:1975:DCS

[Din75b] Nelson T. Dinerstein. Does computer science belong in a liberal arts college? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):55–64, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dinerstein:1977:RSD

[Din77] Nelson T. Dinerstein. A remedial software design course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):77–78, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dyba:1975:CHR

[DKMS75] Jerome Dyba, Marvin Kornbluh, William McCartin, and Richard G. Schneider. The challenge of human resources staffing and utilization in industrial computing — part 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7 (1):123, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Druffel:1977:AWP

[DMH^{+}77] Larry E. Druffel, Vance A. Mall, Robert N. Hawley, Kenneth L. Krause, and Marion A. Pumfrey. Another way to present computer programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):85–87, December 1977. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Danielson:1975:ATI

[DN75] Ronald L. Danielson and Jurg Nievergelt. An automatic tutor for introductory programming students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):47– 50, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Dumas:1976:AC

[DN76] Robert J. Dumas and Anne E. Nieberding. An alternative to Curriculum 68. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):5–11, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Dowling:1974:CGC

[Dow74] Wayne C. Dowling. A computer-graphics course for undergraduate engineers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(2):5–8, June 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Deimel:1979:RSP

[DP79] Lionel E. Deimel, Jr. and Mark Pozefsky. Requirements for student programs in the undergraduate computer science curriculum: How much is enough? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):14–17, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Dexter:1975:DSC

[DRT75] Margaret E. Dexter, Margaret L. Rhoden, and Jerry Sue Townsend. A discrete structures course for a small college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):61–64, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

DuWorks:1972:APD

[DS72] Robert J. DuWorks and Steven W. Smoliar. The arrogant programmer: Dijkstra and Wegner considered harmful. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(4):19–21, December 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DeLorm:1975:MFE

[DS75] R. T. DeLorm and T. C. Smith. Motivating freshmen engineering students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):124–128, February 1975. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Davis:1975:IED

[DTW75] Alan Davis, Michael H. Tindall, and Thomas R. Wilcox. Interactive error diagnostics for an instructional programming system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):168–171, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Dunning:1974:ROS

[Dun74] Peter Dunning. Recommendations for an operating systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):90, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Engel:1979:CCP

[EAAD79] Gerald L. Engel, William F. Atchison, Richard H. Austing, and Chester Davis. From Curriculum ’68 to Curriculum ’78 (panel discussion). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):22, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Engel:1972:EEC

[EB72] Gerald L. Engel and Bruce H. Barnes. The effect of environment on computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):13–18, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eastman:1979:RIC

[EB79] Caroline M. Eastman and Steven L. Blumsack. Repackaging the introductory course to separate the control language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):155–157, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Eckhouse:1972:CSL

[Eck72] Richard H. Eckhouse. The computer science laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):42–45, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eckberg:1976:SPD

[Eck76a] Carl F. Eckberg. Some proposals for distributing central computing power at a university. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):129–134, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Ecklund:1976:LNP

[Eck76b] E. F. Ecklund, Jr. A “non-programming” introduction to programming concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):62–64, February 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Eckberg:1978:QKK

[Eck78] Carl Eckberg. QUILT (a.k.a. KWILT): a special purpose instructional language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):126–130, August 1978. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Ein-Dor:1970:SAC

[EDL70] Phillip Ein-Dor and Norman Lyons. Systems analysis in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(5):16–21, December 1970. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eccles:1976:CST

[EG76] William J. Eccles and Brian G. Gordon. Computer science by TV. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):54–56, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Egan:1976:CGB

[Ega76] L. G. Egan. Closing the “gap” between the university and industry in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(4):19–25, December 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Engel:1973:DSU

[EJ73] G. L. Engel and N. D. Jones. Discrete structures in the undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):56–59, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Ellzey:1973:CRO

[Ell73] Roy S. Ellzey. The college’s role in offering short concentrated courses for industry. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(4):11–13, December 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ellis:1975:PSS

[Ell75] Clarence A. Ellis. A practicum — seminar in structured programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):86–94, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Embley:1978:TDF

[Emb78] David W. Embley. Teaching disciplined Fortran programming via unfort. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(2):34–38, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Engel:1970:CUC

[Eng70a] Gerald L. Engel. Comments on an undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(4):54–55, September/October 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Engel:1970:CSE

[Eng70b] Gerald L. Engel. Computer science education in small colleges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(2):14–30, June 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Engel:1971:IAC

[Eng71a] G. Engel. Input from ACM Curriculum Committee on Computer Science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(4):30–39, December 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Engel:1971:CSI

[Eng71b] G. L. Engel. Computer science instruction in small colleges — an initial report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(2):8–18, June 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Engel:1974:IRR

[Eng74] Gerald L. Engel. Initial report: the revision of “Curriculum 68”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):79–80, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Engel:1976:RCA

[Eng76] Gerald L. Engel. The revision of “Curriculum ’68” (an abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):2–3, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Epstein:1975:PTC

[Eps75] G. Epstein. The pyramid teaching computer structures by computer structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):58–61, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Epley:1978:TSC

[ES78] Donald Epley and Ted Sjoerdsma. A two-semester course sequence in introductory programming using PL/1 — a rationale and overview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):113–119, August 1978. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Ellis:1972:TCD

[EW72] Robert A. Ellis and Donald F. Wann. Teaching computer design using macromodules. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):160–162, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Faiman:1974:STR

[Fai74a] Michael Faiman. Some thoughts on the role of hardware in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):103–105, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fairley:1974:GCB

[Fai74b] Richard E. Fairley. A graduate curriculum in Business-Oriented Computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):37–39, February 1974. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Fairley:1978:TMC

[Fai78] Richard E. Fairley. Toward model curricula in software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):77–79, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Fisher:1978:SHP

[FAT78] James Fisher, John Alvarez, and Robert Taylor. A survey of how practicing programmers keep up-to-date first results including their implications for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):67–72, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Fletcher:1976:PDI

[FBB^{+}76] J. D. Fletcher, Avron Bar, John Seely Brown, Donald Gentner, Ira Goldstein, and Mark Miller. Panel discussion: Intelligent instructional systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):371, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Forsythe:1970:CSM

[FGH^{+}70] G. E. Forsythe, B. A. Galler, J. Hartmanis, A. J. Perlis, and J. F. Traub. Computer science and mathematics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(4):19–29, September/October 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Flon:1978:FCS

[FHSW78] Lawrence Flon, Paul N. Hilfinger, Mary Shaw, and Wm. A. Wulf. A fundamental computer science course that unifies theory and practice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):255–259, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Fisher:1973:SIP

[FHW73] P. Fisher, W. Hankley, and V. Wallentine. Separation of introductory programming and language instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):9–14, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Findler:1970:PDQ

[Fin70] Nicholas V. Findler. Ph.D. qualifying exam questions in artificial intelligence and symbol manipulation at the State University of New York at Buffalo. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(5):43–47, December 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Furugori:1977:FCC

[FJ77] T. Furugori and P. Jalics. First course in computer science, a small survey. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):119–122, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Foutz:1978:ITC

[FJ78] Patricia Bartos Foutz and Martha J. Johnson. The integration of a traditional cooperative education program and a computer science discipline. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):87–91, February 1978. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Fu:1973:CSE

[FK73] Kuan-Chen Fu and Benjamin Koo. Computer science education for an overlooked group — high school teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):51–52, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Friedman:1976:SPC

[FK76] Frank L. Friedman and Elliot B. Koffman. Some pedagogic considerations in teaching elementary programming using structured FORTRAN. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):1–10, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Friedman:1977:TPS

[FK77] Frank L. Friedman and Elliot B. Koffman. Teaching problem solving and structured programming in FORTRAN. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):63–68, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Furuta:1979:EEP

[FK79] Richard Furuta and P. Michael Kemp. Experimental evaluation of programming language features: Implications for introductory programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):18–21, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Fletcher:1976:QQI

[FL76] Sharon Fletcher and Thomas Luce. QUICK: a quasi-interactive computer-based instructional system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):275– 279, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Fletcher:1974:CSS

[Fle74] Dexter Fletcher. Computer science seminars on pedagogical techniques and methods for evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):47, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Fosdick:1978:PIC

[FM78] Howard Fosdick and Karen Mackey. A pragmatic introductory course in COBOL: description and rationale. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10 (2):53–58, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fosdick:1979:CPT

[FM79] Howard Fosdick and Karen Mackey. A course on the pragmatic tools of the programming environment: description and rationale. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(3):11–14, September 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Forsythe:1970:SQE

[For70a] G. Forsythe. Syllabi and qualifying examinations for the Ph.D. in computer science at Stanford University (continued). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(1):19–28, March 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Forsythe:1970:PEP

[For70b] George E. Forsythe. Problems, exams, projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(5):37–43, December 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Forbrich:1973:UCD

[For73] Carl A. Forbrich, Jr. Use computers to demonstrate the differential equations of aerodynamics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):119–120, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Frederick:1976:SAC

[Fre76a] Terry J. Frederick. Self-assessment for the computing professional: Consideration of a natural delivery system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8 (3):4, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Freeman:1976:RSD

[Fre76b] Peter Freeman. Realism, style, and design: Packing it into a constrained course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):150–157, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Freeman:1976:SEE

[Fre76c] Peter Freeman. Software engineering education: Needs and objectives. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):266, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Friedman:1975:ETD

[Fri75] Frank L. Friedman. An experience in teaching disciplined programming at an elementary level. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):38–43, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Friedman:1979:UCL

[FS79] Frank Friedman and Judith A. Stebulis. An undergraduate compiler laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):28–36, February 1979. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Gotterer:1973:CSM

[GB73] Malcolm H. Gotterer and Bruce H. Barnes. The computer science M.S. graduate. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):106–109, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Guha:1977:SCH

[GCS77] R. K. Guha, P. A. Carr, and C. L. Smith. Standards considered helpful. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):74–78, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Gerhart:1975:MTP

[Ger75] Susan L. Gerhart. Methods for teaching program verification. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):172–178, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Gorsline:1973:CSE

[GG73] George W. Gorsline and Duff Green III. Computer science education through a rearview mirror: Experiences with Curriculum 68 at Virginia Polytechnic Institute and State University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):102–105, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Grit:1977:TLB

[GG77] Dale H. Grit and Dennis D. Georg. A top-down, laboratory based operating system course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):73–76, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Gruener:1978:SFC

[GG78] William B. Gruener and Steven M. Graziano. A study of the first course in computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):100–107, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Germano:1978:SUH

[GH78] Frank Germano and Stephen Higgenbotham. A student-use: hierarchical data base management system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):53– 58, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Grout:1979:ECS

[GH79] Jarrell C. Grout and Denis M. Hyams. Experiences with a Computer Science Advisory Council. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):127–130, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Gibbs:1977:ICS

[Gib77] Norman E. Gibbs. An introductory computer science course for all majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):34–38, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Gillett:1976:IPA

[Gil76] Will Gillett. An interactive program advising system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):335–341, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Ginsberg:1975:GLU

[Gin75] Myron Ginsberg. A guide to the literature for undergraduate and graduate courses in numerical mathematics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(4):53–68, December 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gries:1974:PDS

[GKC^{+}74] David Gries, Priar W. Kerrighan, Donald R. Chand, John Woolley, Richard C. Holt, Thomas W. S. Plum, Victor Pasill, Olivier Lecarme, and Clirtor R. Foulk. Panel discussion on structured programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):60–68, September 1974. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Glasser:1975:TOH

[Gla75] A. L. Glasser. A terminal oriented hardware simulator for educational use. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):179–186, February 1975. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Givbbs:1974:HCS

[GLO74] Norman E. Givbbs, Donald W. Loveland, and James M. Ortega. The heart of computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(4):13–14, December 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Garland:1979:NMD

[GM79] Stephen J. Garland and A. Kent Morton. A new Master’s degree program at Dartmouth College. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(4):25–26, December 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goddard:1973:SCS

[God73] Alton R. Goddard. Structure and content of service courses in computer science for other disciplines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):15– 17, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Goddard:1974:RCS

[God74] Alton R. Goddard. Relating computer science programs to the needs of industry through internships and cooperative programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):165–168, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Goddard:1975:CAP

[God75] Alton R. Goddard. Computer applications for prospective public school administrators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):51–55, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Golde:1971:ISC

[Gol71] Hellmut Golde. Introducing students to computer science literature — an approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(2):21–25, June 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gonzales:1978:CED

[Gon78] Carlos Gonzales. A computer engineering degree in Mexico. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):48–52, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Graham:1970:TSP

[Gra70] Robert M. Graham. Teaching systems programming and software design: problems and solutions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):56– 60, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grant:1978:ICT

[Gra78] John Grant. Individualized course in theoretical computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(2):39–41, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gries:1974:WSW

[Gri74] David Gries. What should we teach in an introductory programming course? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):81–89, February 1974. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Grillo:1978:SSU

[Gri78] John P. Grillo. Small systems for undergraduate courses in system design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):282–285, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Gross:1975:VAC

[Gro75] James F. Gross. Video augmented computer science (VACS). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(4):47–49, December 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Giaccai:1975:ES

[GS75] Gerald J. Giaccai and Kenneth R. Slonneger. Enhancing SAMOS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):73–76, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Golde:1976:WSC

[GS76] Hellmut Golde and Alan Shaw. Why a separate computer facility for computer science education? experiences at the UW computer sciences teaching laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(2):11–20, June 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gumb:1977:LEE

[Gum77] Raymond D. Gumb. Language extensibility in extended Fortran: data types and pointer variables. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):53–54, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gyllenskog:1978:TOS

[Gyl78] Joel Gyllenskog. Teaching operating systems design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(2):44–46, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Heilman:1971:RED

[HA71] R. L. Heilman and G. P. Ashby. Re-evaluation of debugging in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(4):15–18, December 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hatfield:1978:CEM

[HA78] Frank J. Hatfield and Roberta R. Alexander. Cooperative education: making it work. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):92–95, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Haag:1978:CUS

[Haa78] James N. Haag. Careers of University of San Francisco computer science graduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):59–62, August 1978. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Hammer:1976:RCS

[HAB^{+}76] Preston Hammer, Richard Austing, Bruce Barnes, Gerald Engel, and Gordon Stokes. Recommendations on computer science curriculum for undergraduate degree programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8 (1):95, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Hager:1976:EDP

[Hag76] Oberita Hager. An evaluation of the data processing program at Eastern Kentucky University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):307– 312, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Hammer:1970:UCS

[Ham70] Preston C. Hammer. Undergraduate computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):1–5, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamblen:1974:SCS

[Ham74] John W. Hamblen. Statistics for computer scientists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):109–111, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamblen:1975:CAG

[Ham75a] John W. Hamblen. The case for an advanced graduate record examination in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):1–6, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Hamblen:1975:CMU

[Ham75b] John W. Hamblen. Computer manpower in the United States — supply and demand. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(4):25–43, December 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamming:1975:PCS

[Ham75c] R. W. Hamming. A philosophy for computer science or my prejudices and confusions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(4):16–18, December 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamblen:1977:CMU

[Ham77] John W. Hamblen. Computer manpower in the United States — supply and demand: an update. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):19–37, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamblen:1979:MPC

[Ham79] John W. Hamblen. Masters program in computer science (panel discussion): Preliminary recommendations of the committee on the curriculum in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):60, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Hansen:1978:RUM

[Han78] William A. Hansen. A report on the use of multimedia courses in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):35, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Harris:1974:RCS

[Har74] Fred H. Harris. The role of computer science education in certification of computer professionals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):1–5, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Hardell:1977:NBL

[Har77] Mary M. Hardell. The near blind leading the blind: One partial solution to the hacker battered student syndrome. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9 (3):63–65, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Harrow:1978:HSS

[Har78] Keith Harrow. How to show something is not: Proofs in formal language and computability theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):27–30, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Hays:1974:UEC

[Hay74] Bill R. Hays. Undergraduate education in computer science at brigham Young university. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):21–27, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haskell:1976:USF

[HBJ76] R. E. Haskell, D. E. Boddy, and G. A. Jackson. Use of structured flowcharts in the undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):67–74, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Hedrick:1972:PMT

[Hed72] G. E. Hedrick. A pedagogical model for top-down syntax analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(2):11–13, July 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hedrick:1977:AIO

[Hed77] G. E. Hedrick. ALGOL68 instruction at Oklahoma State University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):16–20, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Helkanoff:1973:MPC

[Hel73] M. A. Helkanoff. An M.S. program in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):77–82, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Holland:1978:HLW

[HFL^{+}78] Robert Holland, David Fulton, Ed Lamie, George Nagy, John Tartar, and Jay Wienkam. Hardware laboratories: what they should contain and how to obtain funding. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):178, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Hyde:1979:IPS

[HGU79] Daniel C. Hyde, Ben D. Gay, and Donald Utter, Jr. The integration of a Problem Solving Process in the first course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):54–59, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Horowitz:1973:CSI

[HH73] E. Horowitz and M. C. Horowitz. Computers and society: an interdisciplinary approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):134–137, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Hansen:1978:TDU

[HHD78] John C. Hansen, Elizabeth A. Hansen, and Kenneth H. Derus. Team dynamics and the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):28–29, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haynes:1977:OTL

[HHP77] William R. Haynes, Jr., Charles E. Hughes, and Charles P. Pfleeger. ORACLE a tool for learning compiler writing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):37–52, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Heimer:1973:TTC

[HJ73] Ralph T. Heimer and Lars C. Jansson. Teacher training in computer education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):48–50, February 1973. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Hamblen:1979:SDP

[HLEA79] John W. Hamblen, Joyce C. Little, Gerald L. Engel, and Richard E. Austing. Status of degree programs in computer science (panel discussion). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):1, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Hanson:1975:FCC

[HM75] Allan Hanson and Kurt Maly. A first course in computer science: What it should be and why. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):95–101, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Hobson:1978:CSH

[Hob78] Richard F. Hobson. Computing science hardware laboratories and the LSI revolution. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):160–170, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Hoffman:1974:COC

[Hof74] Lance J. Hoffman. Course outline for computer security and privacy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):13–17, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hoffman:1978:SSE

[Hof78] A. A. J. Hoffman. A survey of software engineering courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):80–83, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Hoffman:1979:MSE

[Hof79] A. A. J. Hoffman. Master of software engineering — a proposed curriculum for practitioners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):70–74, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Homeyer:1977:EMC

[Hom77] Fred C. Homeyer. An experimental microcomputer course (a case history). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):41–44, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Homeyer:1978:SGU

[Hom78] Fred C. Homeyer. A survey of graduates from an undergraduate computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):31, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Honkanen:1973:SCP

[Hon73] Pentti A. Honkanen. A student-computer programming system for teaching graph and network theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):177– 179, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Honkanen:1975:SFC

[Hon75] Pentti A. Honkanen. Setting the foundations of computer science in a business oriented program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):187–190, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Hosch:1973:SCR

[Hos73] Frederick A. Hosch. Some comments on the role of computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(3):13–17, September 1973. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hosch:1977:WF

[Hos77] Frederick A. Hosch. Whither flowcharting? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9 (3):66–73, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Hughes:1976:AVT

[HP76] Charles E. Hughes and Charles P. Pfleeger. ASSIST-V: a tool for studying the implementation of operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):167–173, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE– SIGCUE joint symposium on Computer science education.

Hilfinger:1978:ITS

[HSW78] Paul N. Hilfinger, Mary Shaw, and Wm. A. Wulf. Introducing “theory” in the second programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10 (3):55–58, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Highland:1973:ITF

[HT73] Harold Joseph Highland and Charles B. Thompson. An integrated [two- and four- year] curriculum in information processing at a state technical college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):86–90, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Huggins:1970:UFP

[Hug70] Phyllis Huggins. Universities failing to provide adequate background for D.P. [report on ACM SIGCSE, FJCC 1970]. ComputerWorld, November 25, 1970. CODEN CMPWAB. ISSN 0010-

4841. URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Compiler/sigada.bib.

Hughes:1977:TDF

[Hug77] Herman D. Hughes. A tool designed to facilitate structured programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):26–30, February 1977. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Hunt:1970:CST

[Hun70] Earl Hunt. The computer science teaching laboratory at the University of Washington. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):30–33, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holt:1974:SSS

[HW74] Richard C. Holt and David B. Wortman. A sequence of structured subsets of PL/I. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):129–132, February 1974.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Hasenei:1975:CHR

[HWC75] Ken Hasenei, Claude Walston, and Ed Coady. The challenge of human resources staffing and utilization in industrial computing — part 2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):157, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Ivey:1972:PLT

[IE72] William Max Ivey and Larry C. Eversole. A programming language for the teaching of algorithmic analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):136– 140, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Isaacs:1976:BCA

[IN76] Gerald L. Isaacs and Melvin R. Novick. The Bayesian computerassisted data analysis (CADA) monitor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):268– 274, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Ingram:1975:RSI

[Ing75] G. Ingram. Report of special interest session on computing in minority institution held at fifth annual SIGCSE symposium. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):12, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Irby:1977:TSD

[Irb77] Thomas C. Irby. Teaching software development using a microprocessor laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):113–118, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Iyengar:1976:EDS

[Iye76] S. Sitharama Iyengar. An experience with a “Data Structures and File Management” course at the graduate level in computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(4):12–14, December 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Iyengar:1978:ECP

[Iye78] S. Sitharama Iyengar. An experience with a computers and programming course at the graduate level in a computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):35–36, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jackson:1976:SET

[Jac76a] Alyce Jackson. In-service education for teachers of computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):29–34, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Jackson:1976:GTD

[Jac76b] Glenn A. Jackson. A graphical technique for describing recursion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):30–32, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Jalics:1978:NTS

[Jal78] Paul J. Jalics. A note on teaching structured programming in an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):79, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Jarvinen:1977:NEP

[Jär77] Pertti Järvinen. Notes on educational planning: a systems approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):57–62, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jehn:1976:ACS

[JBF76] Lawrence A. Jehn, James E. Brandeberry, and B. Albert Friedman. Articulation of computer science curricula at the community colleges and the universities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):149, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Jehn:1974:JSC

[Jeh74] L. A. Jehn. Joint session computer science and engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(4):21–25, December 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:1977:UCS

[JK77] Warren T. Jones and Donald L. Kalmey. The utility of computer simulation as instructional devices for operating systems courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):55–56, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:1975:NCS

[Jon75a] Clinton E. Jones. Need for computer-supported instruction in minority institutions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):12–20, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:1975:POC

[Jon75b] Warren T. Jones. A pragmatically oriented computer science degree program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):65–66, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:1979:TAD

[Jon79] Bush Jones. Teaching algorithm design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(4): 27–30, December 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jordan:1972:RCS

[Jor72] Sara R. Jordan. Review of “Computer Science Laboratory Exercises” by F. D. Federighi and E. D. Reilly, Jr. Riedinger & Riedinger, Schenectady, N. Y., 1971. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(3):36, October 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jehn:1978:CSE

[JRS78] Lawrence A. Jehn, David C. Rine, and Norman Sondak. Computer science and engineering education: Current trends, new dimensions and related professional programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):162–178, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Jensen:1978:PYS

[JTF78] Randall W. Jensen, Charles C. Tonies, and William I. Fletcher. A proposed 4-year software engineering curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10 (3):84–92, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Johnston:1973:PPF

[JW73] August L. Johnston, Jr. and Theodore C. Willoughby. Programming from prose, flowcharts or decision tables. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(4):4–7, December 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:1978:CPV

[JW78] Gerald A. Jones and Ann Marie Walsh. A course in program verification for programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):213–216, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Kalmey:1974:PCI

[Kal74] Donald L. Kalmey. Profile of a computer and information science B.S. graduate. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):40–45, February 1974. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Kay:1974:CSR

[Kay74] Alan C. Kay. Computer scientists’ responsibility to educational research: a dynamic medium for creative thought. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6 (1):46, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Kearney:1975:CTY

[Kea75] James F. Kearney. Curricula for two-year data-processing programs: a review and recommendation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):143– 147, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Keith:1976:SBC

[Kei76] Ed Keith. A structured beginning COBOL class using structured programming with objectives. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):143–147, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Kerr:1978:PAI

[Ker78] Douglas S. Kerr. Programming and algorithms II: a second course in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):45–49, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Koffman:1976:CAF

[KF76] Elliot B. Koffman and Frank L. Friedman. A computer-aided flow diagram teaching system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):350–354, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Konstam:1974:ALF

[KH74] Aaron H. Konstam and John E. Howland. APL as a lingua franca in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):21–27, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Khallany:1975:ICC

[KH75] Asad Khallany and Robert H. Holland. An introductory computer course in a school of business. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):39–42, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Khailany:1976:ICC

[Kha76] Asad Khailany. An introductory COBOL course with structured programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):11–16, February 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Khailany:1977:ASC

[Kha77a] Asad Khailany. Advanced structured COBOL programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):59–62, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Khailany:1977:ATS

[Kha77b] Asad Khailany. Alternative teaching strategy for an introductory computer language course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):93–95, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Kline:1978:VDP

[KHKD78] Russell B. Kline, Gary D. Hamor, Kenneth L. Krause, and Larry E. Druffel. Visual demonstration of program execution. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):16–18, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/ CSA Technical Symposium on Computer Science Education.

Kimura:1977:RPE

[Kim77] Takayuki Kimura. Recursive programming in English for freshmen. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):129–132, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Kimura:1978:HGC

[Kim78] Izumi Kimura. A historical, generalistic, and complementary approach in introductory computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):93–99, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Kimura:1979:RBC

[Kim79] Takayuki Kimura. Reading before composition. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):162–166, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Kerr:1978:BMC

[KK78] Douglas S. Kerr and Donald L. Kalmey. The Bachelor’s and Master’s computer science graduate. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):36, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Khalil:1978:AIC

[KL78] Hatem Khalil and Leon S. Levy. The academic image of computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(2):31–33, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koltun:1979:CLC

[Kol79] Philip Koltun. “computers and the law”: a course description. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):236–239, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Komar:1970:LAE

[Kom70] Joseph A. Komar. Liberal arts education in computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):84–105, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Konstam:1975:SPP

[Kon75] Aaron H. Konstam. Structured programming and the parallel algorithm. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):43–47, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Korfhage:1970:VIP

[Kor70] Robert R. Korfhage. Visiting industrial professors and sabbaticals in industry. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):9–12, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Korfhage:1975:CBR

[Kor75] Robert R. Korfhage. The case for B3 — a reply to Waxman. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(4):69–71, December 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). See [Wax75].

Kernighan:1974:PS

[KP74] B. W. Kernighan and P. J. Plauger. Programming style. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):90–96, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/

att.bib. Proceedings of the 4th SIGCSE symposium on Computer science education.

Kennedy:1976:EEL

[KP76] William E. Kennedy and James R. Pinkert. EDUSAC — an educational list processing system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):33–38, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Kuo:1974:AHC

[KPM74] M. H. Kuo, G. A. Perdikaris, and T. E. Mueller. An analog/hybrid computer simulation of phase-locked-loops in motor speed control. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):101–105, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Kildall:1972:AEA

[KR72] Gary A. Kildall and Alan B. Roberts. ALGOL-E: an experimental approach to the study of programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):127–135, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:1976:IUO

[KR76] V. K. Kumar and James L. Rogers. Instructional uses of the olin experimental classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):189–191, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Kenworthy:1979:STP

[KR79] D. J. Kenworthy and K. A. Redish. Software team projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):37–40, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Krall:1972:RUR

[Kra72] Edward J. Krall. Results of the use of a recursive function translator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):110–115, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

King:1979:CCI

[KRFS79] Ronald S. King, David C. Rine, DuWayne D. Furman, and William A. Schaefer. Competency in computer instruction in Illinois Public Schools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):89–100, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Krueger:1978:RTL

[Kru78] Myron W. Krueger. Real-time laboratory for interdisciplinary computer projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):286–290, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Korfhage:1974:IIC

[KS74] Robert R. Korfhage and Robert J. Smith. Individualized instruction in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):161–164, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Kirbach:1976:UME

[KS76] Jürgen Kirbach and Joachim W. Schmidt. On the uses of models in education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):212–222, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Khailany:1978:CPT

[KS78] Asad Khailany and Charles Saxon. Conducting project team classes in data processing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):189–192, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Kugel:1975:HMA

[Kug75] Peter Kugel. How to make abstract ideas more concrete. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):191–195, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

King:1978:CPA

[KW78] Ronald King and John Walstrom. A course in personalized applications libraries. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):73–76, August 1978. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Lake:1976:TPC

[Lak76] Robin B. Lake. A training program in computer applications in the health sciences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):89–91, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Lamie:1976:UGT

[Lam76] Edward L. Lamie. Using GPSS to teach operating systems concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):174–178, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Lamie:1978:UID

[Lam78] Edward L. Lamie. Using an instructional data base system in the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):59–64, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Lane:1975:HAT

[Lan75] Malcolm G. Lane. A hands-on approach to teaching systems programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):23–30, February 1975. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Little:1977:CRG

[LAS^{+}77] Joyce Currie Little, Richard H. Austing, Harice Seeds, John Maniotes, and Gerald L. Engel. Curriculum recommendations and guidelines for the community and junior college career program in computer programming: a working paper of the association for computing machinery committee on curriculum in computer sciences by the sub committee on community and junior college curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(2):17–36, June 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lavington:1978:PFC

[Lav78] S. H. Lavington. Problems faced by a computer science department in a developing country. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):128–131, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Lawes:1974:DIC

[Law74] C. Peter Lawes. Description of an introductory course in programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):69–78, September 1974. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lecarme:1974:SPP

[Lec74] Olivier Lecarme. Structured programming, programming teaching and the language Pascal. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(2):9–15, June 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:1972:CSC

[Lee72] Hans E. Lee. Computers in society: a course description, purpose and rationale. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):97–102, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leininger:1972:CRS

[Lei72] C. W. Leininger. Computer related studies at a college of arts, sciences and education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(3):18–35, October 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leinbaugh:1977:DPS

[Lei77] Dennis Leinbaugh. A demand paging simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9 (3):43–48, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Lemos:1978:CET

[Lem78a] Ronald S. Lemos. The cost-effectiveness of team debugging in teaching Cobol programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):193–196, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Lemos:1978:FVF

[Lem78b] Ronald S. Lemos. Factionalism versus fraternalism in computing: a plea for the latter. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):16–22, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Lemos:1979:TPL

[Lem79] Ronald S. Lemos. Teaching programming languages: a survey of approaches. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):174–181, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Levine:1976:CCD

[Lev76] David R. Levine. Computer-controlled display demonstrations of dynamic concepts in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):192– 199, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Levine:1977:SIP

[Lev77] David R. Levine. The simple internal procedure as an effective tool in the teaching and practice of structured PL/I. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):22–25, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Levine:1978:WIP

[Lev78] David R. Levine. A workable interpreter project for a “Structured Programming” course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):38–44, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Levin:1979:USB

[Lev79] Peter Levin. Using a single board microcomputer with a display in a computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):144–148, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Lewis:1974:MAI

[Lew74] T. G. Lewis. Minicomputer assisted instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6 (3):46–50, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lewis:1975:ICS

[Lew75] Jesse C. Lewis. Innovative computer services for minority colleges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):7–10, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Lewis:1976:CSM

[Lew76] Jesse Lewis. Computer science and minority colleges & universities: The Jackson state university NSF educational computing network. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):382–383, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Lee:1979:DEC

[LHL79] Iva Helen Lee, Donna Hutcheson, and Joyce Currie Little. Data entry curricula guidelines: a working paper of the Community and Junior College Subcommittee (CAJC) of the Curriculum Committee on Computer Education (C^{3}E). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(3):15–39, September 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lewis:1976:UGP

[LI76] Jesse C. Lewis and S. Sitharama Iynengar. A unique graduate program in computer science at Jackson state university. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):355–358, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE– SIGCUE joint symposium on Computer science education.

Linder:1976:CTS

[Lin76] William H. Linder. COMPUTER-TUTOR: From a student project to a self-paced CAI/CMI course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):57–60, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Lin:1978:MCS

[Lin78] Benjamin M. Lin. Microprocessors in a computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):31–34, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Linder:1979:CSM

[Lin79] William H. Linder. The computer science minor, a description and a proposal. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(2):40–42, June 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lions:1970:PEP

[Lio70] John Lions. Problems, exams, projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(2):37– 41, June 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Little:1974:CCC

[Lit74] Joyce Currie Little. The community colleges and Curriculum ’68. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):98–102, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Little:1977:RTT

[Lit77] Joyce C. Little. Report on three tutorials for SIGCSE Bulletin proceedings. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):7, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lorton:1976:MSD

[LK76] Paul Lorton, Jr. and Rosemary N. Killam. Modeling the student and the discipline in CAI drill and practice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):223–235, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE– SIGCUE joint symposium on Computer science education.

Leither:1978:WJC

[LL78] Henry E. Leither and Harry R. Lewis. Why johny can’t program: a progress report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):266–276, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Lowther:1976:TGP

[LM76] J. L. Lowther and Z. C. Motteler. Teaching good programming techniques. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(4):10–11, December 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lopez:1978:IAC

[Lop78] Antonio M. Lopez, Jr. An implementation of ACM Curriculum 77. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(2):47–52, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Loveman:1972:CAP

[Lov72] David B. Loveman. A course in advanced programming for undergraduate computer science majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):27–36, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Loveland:1974:PSM

[Lov74] D. Loveland. Position statement: mathematical computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):106–108, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lovas:1975:ICP

[Lov75] Charles M. Lovas. An instructional computer program library. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):129–132, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Lovas:1979:TCA

[Lov79] Charles M. Lovas. Teaching computer applications using computer cases. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(2):23–26, June 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

LaFrance:1972:CSL

[LR72] Jacques LaFrance and Waldo Roth. Computer science for liberal arts colleges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(4):22–31, December 1972. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

LaFrance:1973:CSL

[LR73] Jacques LaFrance and R. Waldo Roth. Computer science for liberal arts colleges: a report of a workshop held at Wheaton College, Wheaton, Illinois, July 12–14, 1972. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):70–76, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Little:1976:RCR

[LSA^{+}76] Joyce Currie Little, Bobby Smith, Richard Austing, Eric Whiteside, and Charles Leidlich. A report on the curriculum recommendations of the ACM Sub-Committee for Community and Junior College Curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):93, July 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Little:1976:RAA

[LSL^{+}76] Joyce Currie Little, Harice Seeds, Ronald Lenhardt, John Dineen, and John Maniotes. Report on ACM’s Activity on Community and Junior College Curriculum Career Program in Computer Programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):267, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Lancaster:1974:ECQ

[LT74] Ronald L. Lancaster and Richard T. Thomas. Experience with Curriculum ’68 and the quarter system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(2):16– 19, June 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lancaster:1975:CSP

[LT75] Ronald L. Lancaster and Richard T. Thomas. A computer science practicum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):76–78, February 1975. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Lancaster:1976:RF

[LT76] Ronald L. Lancaster and Richard T. Thomas. Reading Fortran. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(4):6–7, December 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lochovsky:1978:TDM

[LT78] F. H. Lochovsky and D. C. Tsichritzis. Teaching data management using an educational data base management system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):65–69, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/ CSA Technical Symposium on Computer Science Education.

Lucas:1976:PAC

[Luc76] William R. Lucas. Planned attitude change while teaching computer literacy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):90–94, February 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Lusk:1977:AUC

[Lus77] Ewing L. Lusk. An advanced undergraduate course in applied computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):28–29, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

MacEwen:1974:PPC

[Mac74] Glenn H. MacEwen. A programming project for a course in operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(2):29–36, June 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

MacEachern:1979:SPA

[Mac79] Alexander MacEachern. Sample programming assignment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(4):7–10, December 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Malkin:1972:CST

[Mal72] Judith G. Malkin. Computer science for teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4 (1):60–66, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Maniotes:1974:SUC

[Man74] John Maniotes. The state of undergraduate computer and data processing programs at public universities in indiana. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):53–58, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Mapp:1973:PBI

[Map73] George A. Mapp. A proposal for a B.S. in information systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):91–94, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Martin:1972:CAI

[Mar72] Don Martin. Computer aided instruction in system dynamics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):53–56, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marshall:1975:VU

[Mar75] Sister Patricia Marshall. The view from down under. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):11–14, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Marsh:1976:TTA

[Mar76] Barbara Marsh. Teaching teachers about computers: a course description. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):86–89, February 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Mashey:1973:ATY

[Mas73] John R. Mashey. ASSIST: Three year’s experience with a studentoriented assembler. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):157–165, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Matula:1970:CPD

[Mat70a] D. W. Matula. Course and program descriptions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2 (1):12–15, March 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Matula:1970:ECA

[Mat70b] David W. Matula. The emergence of computational arithmetic as a component of the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):41–44, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mathis:1974:TD

[Mat74] Robert F. Mathis. Teaching debugging. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):59– 63, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Mathis:1978:FCC

[Mat78] Robert F. Mathis. A first course in computer science for graduate students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):217–223, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Maurer:1971:PDP

[Mau71] Ward Douglas Maurer. Ph. D prelim exams in computer systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(1):46–47, March 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Maurer:1977:TPC

[Mau77] W. D. Maurer. The teaching of program correctness. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):142–144, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Mavaddat:1976:ETP

[Mav76] F. Mavaddat. An experiment in teaching programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(2):45–59, June 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mayes:1975:DCF

[May75] Jesse J. Mayes. Development of computing at Federal City College. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):114–119, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Mazlack:1976:DCS

[Maz76] Lawrence J. Mazlack. Does a computer have sexual preferences? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):74–78, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Mazlack:1977:DCA

[Maz77] Lawrence J. Mazlack. Developing computer awareness. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):184–187, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Mazlack:1978:UIF

[Maz78] Lawrence J. Mazlack. The use of interactive Fortran. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):260–265, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/ CSA Technical Symposium on Computer Science Education.

Mazlack:1979:RCS

[Maz79] Lawrence Mazlack. The role of computer science education in aiding technology transfer to less developed countries (panel discussion). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):214, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Meinke:1976:IMC

[MB76a] John G. Meinke and Charles R. Bauer. The IIT MST in computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):35–38, February 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Meinke:1976:TST

[MB76b] John G. Meinke and Charles R. Bauer. Topological sorting as a tool in curriculum planning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):61–66, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Martin:1977:PFP

[MB77] Edith W. Martin and Albert N. Badre. Problem formulation for programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):133–138, February 1977. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Meinke:1979:SES

[MB79] John G. Meinke and John A. Beidler. A software emphasis in the second course in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):192–194, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

McCharen:1974:DFC

[McC74] John D. McCharen. A description of a first course in assembler language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):36–43, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCurdy:1978:TTE

[McC78] Robert McCurdy. Teaching toward error free programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):33, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/ CSA Technical Symposium on Computer Science Education.

McFarland:1975:NPD

[McF75] Robert H. McFarland. Notes from panel discussion on GRE advanced exams presented at fifth annual SIGCSE technical symposium. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(4):44–46, December 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McGinley:1973:TTU

[McG73] Pamela R. E. McGinley. The training of teachers in the use of computers in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):53–55, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

McAdams:1976:CGA

[MD76] Joseph K. McAdams and Arlan R. DeKock. Computer graphics as an aid to teaching geometric transformations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):137–143, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE– SIGCUE joint symposium on Computer science education.

Mein:1977:TSP

[Mei77] William J. Mein. Toward the specification of programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):139–141, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Mein:1979:CAC

[Mei79] Wm J. Mein. Computer arts: Computer science in a liberal arts environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):224–227, February 1979. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Menninga:1974:IPE

[Men74] Larry D. Menninga. Introducing practical experience into Curriculum 68 through integration of courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):152– 154, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Metzner:1979:CPD

[Met79] John R. Metzner. Contesting (panel discussion). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11 (1):75, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Mackey:1979:ACS

[MF79] Karen Mackey and Howard Fosdick. An applied computer science/systems programming approach to teaching data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):76–78, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Meissner:1975:BSM

[MH75] Loren P. Meissner and Ruth L. Hinkins. B4Tran: a structured mini-language approach to the teaching of Fortran. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):200–205, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Moshell:1978:ISC

[MHA78] J. M. Moshell, C. E. Hughes, and R. M. Aiken. An integrated small computer laboratory using a central disk storage facility. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):35–37, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Michalopoulos:1976:VDO

[Mic76] Demetrios A. Michalopoulos. A video disc oriented educational system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):389–392, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Mijares:1978:GPI

[Mij78] C. Ignacio Mijares. A graduate program in information systems for the Latin American environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):132–136, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Mills:1973:SCP

[Mil73] H. D. Mills. On the structure of a computing profession. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):97–101, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Miller:1975:IOC

[Mil75] Jim Miller. An industry oriented computer curriculum designed for the transfer student. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(4):50–52, December 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Miller:1978:IIM

[Mil78] James E. Miller. Iterations on an instructional method for utilizing the computer in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):30–31, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mitchell:1979:AAS

[Mit79] William Mitchell. Another approach to service courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):6–9, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Mize:1976:MAC

[Miz76a] J. L. Mize. Making an academic curriculum relevant to business requirements. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(2):24–27, June 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mize:1976:SPC

[Miz76b] Jan L. Mize. Structured programming concepts in an introductory COBOL course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):135–142, July 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Mathis:1972:DMU

[MK72] Robert F. Mathis and Douglas S. Kerr. Development of a multifacetted undergraduate program in computer and information science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):8–12, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morgan:1973:ALP

[MK73] Howard L. Morgan and James C. Kinard. ASAP: a language and philosophy for teaching file processing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):21– 23, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Moshos:1974:FRP

[MK74] George J. Moshos and Kevin B. Kern. Final report of Phase III Master Plan Committee for Computer Science and Technology to New Jersey Department of Higher Education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6 (2):24–28, June 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:1973:SHA

[MM73] Johannes J. Martin and Therese Martin. Student help and automatic grading: The organization of a large service course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):138–141, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Magel:1978:PMD

[MM78a] Kenneth Magel and Everald E. Mills. A professional Master’s degree program in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):27–30, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Mitchell:1978:ICS

[MM78b] William Mitchell and Bruce Mabis. Implementing a computer science curriculum merging two curriculum models. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):151–155, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Moursund:1976:CSE

[MN76] David Moursund and Mike Neill. Computer science for elementary school teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):24–28, February 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Mocciola:1978:TTC

[Moc78] Michael R. Mocciola. Teacher-training in computer science education in Western Australia: group projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):206–209, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Modesitt:1973:CSU

[Mod73] Kenneth L. Modesitt. Computer science at a university regional campus. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(2):21–22, June 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Modesitt:1974:PVA

[Mod74] Kenneth L. Modesitt. PSI: a valuable addition to the alphabet soup for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(2):37–44, June 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moon:1970:CCS

[Moo70] B. A. M. Moon. The challenge of computer science in New Zealand. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):61–68, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moore:1979:FCF

[Moo79] Freeman L. Moore. A first course on files. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):79– 81, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Morris:1972:CSS

[Mor72] Glenn Morris. Computer supported study of chemical equilibria. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):141–143, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morton:1976:VTS

[Mor76] Richard P. Morton. The variety of TICCIT systems — an overview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):144–148, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Morrison:1977:TCF

[Mor77] Louise S. Morrison. Training college faculty members in the educational uses of computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):173–174, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Moursund:1974:CPC

[Mou74] David Moursund. Computers in pre-college education: Oregon moves forward. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):169–173, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Moursund:1976:CCS

[Mou76] David Moursund. Calculators and the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(2):21–23, June 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moursund:1977:CMG

[Mou77] David Moursund. Calculator metaphors, and goals for calculator education in elementary schools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):100– 103, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Moursund:1978:OCS

[Mou78] David Moursund. Opinions on computer science courses for precollege level teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):183–188, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Moursund:1979:CPE

[Mou79] David Moursund. Computers in precollege education (panel discussion): What do teachers want or need? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):140, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Murray:1978:UCS

[MPS78] Bob Murray, Tom Policano, and John Sweeney. Using captioned slides to teach computer operations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):80, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

McDaneld:1976:ITP

[MS76a] T. G. McDaneld and E. J. Schweppe. An interactive turning/post/ mixed machine simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(2):36–40, June 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morrison:1976:ACA

[MS76b] Louise S. Morrison and Grover C. Simmons. Academic computing at Atlanta University Center — a consortium of six predominantly black institutions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):95–97, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Maguire:1977:ELF

[MS77] R. Brien Maguire and L. R. Symes. Effect of laboratory facilities on computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):1–5, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Marsland:1973:CMS

[MT73] T. A. Marsland and J. Tartar. A course in minicomputer systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):153–156, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Mullins:1976:CEC

[Mul76] E. R. Mullins, Jr. Computer education for computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(4):4–5, December 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Muscat:1976:SSM

[Mus76] Eugene J. Muscat. Secondary school model automatic test generation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):113–115, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Nahapetian:1979:CSS

[Nah79] Armen Nahapetian. Computer science (subjectively) defined. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(4):11–14, December 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nanney:1976:CSE

[Nan76] T. Ray Nanney. Computer science: an essential course for the liberal arts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):102–105, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Nartker:1970:CPD

[Nar70] T. A. Nartker. Course and program descriptions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2 (4):41–53, September/October 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Neidleman:1976:MSP

[Nei76] L. D. Neidleman. The Master of Science Program in Computer and Information Science at San Jose State University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):359–365, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE– SIGCUE joint symposium on Computer science education.

Neill:1977:STR

[Nei77] Michael J. Neill. Some thoughts on reasons, definitions and tasks to achieve “functional” computer literacy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):175– 177, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Nelson:1971:KSU

[Nel71] Bradley Nelson. Kansas State University’s traveling seminar. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(2):19–21, June 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nelson:1978:AIT

[Nel78] Randolph Nelson. An approach to improved textbook writing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):47–51, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Newman:1973:ATT

[New73] J. Richard Newman. Alternative teaching techniques in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(4):29–32, December 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Newsted:1975:GAP

[New75] Peter R. Newsted. Grade and ability predictions in an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):87–91, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nievergelt:1978:XSE

[NFB^{+}78] J. Nievergelt, H. P. Frei, H. Burkhart, C. Jacobi, B. Pattner, H. Sugaya, B. Weibel, and J. Weydert. XS-0: a self-explanatory school computer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):66–69, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nielsen:1970:NCC

[Nie70a] Norman R. Nielsen. Network computing for computer science departments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):24–29, November 1970. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nievergelt:1970:SLL

[Nie70b] J. Nievergelt. Syllabus for low level performance for the qualifying examination in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(2):43–53, June 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nielsen:1972:SRC

[Nie72] Norman R. Nielsen. Social responsibility and computer education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):90–96, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Noe:1973:UEC

[Noe73] Jerre D. Noe. University education in computer measurement and evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(2):8–12, June 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Noonan:1979:SCC

[Noo79] Robert E. Noonan. The second course in computer programming: Some principles and consequences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):187–191, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Norris:1974:ADP

[Nor74] Fletcher R. Norris. An algorithm for determining perfect squares. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):8–12, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nievergelt:1973:AIC

[NR73] Jurg Nievergelt and Edward M. Reingold. Automating introductory computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):24–25, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Newton:1976:TBP

[NS76] Glen E. Newton and J. Denbigh Starkey. Teaching both PL/I and Fortran to beginners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):106–107, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Nievergelt:1970:CES

[NT70] J. Nievergelt and Wenneth J. Travers. Computer education for secondary school mathematics teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):75– 83, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nievergelt:1974:TMA

[NW74] Jurg Nievergelt and Thomas R. Wilcox. From teaching-machine to automatic tutor: The changing man-machine interface. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):46, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Nagy:1978:PPP

[NW78] G. Nagy and L. Wilson. Program portability from the perspective of students at the University of Nebraska. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(2):59– 62, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Oliver:1973:NUC

[Oli73] James R. Oliver. The need to upgrade computer science curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(4):14–18, December 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Olivieri:1974:PLU

[Oli74] C. Peter Olivieri. Programming language usage and areas of concentration in computer sciences education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6 (3):33–35, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Omlor:1978:SPC

[Oml78] J. Dennis Omlor. Structured programming for computer science majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):34, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Oldehoeft:1977:MTI

[OR77] R. R. Oldehoeft and R. V. Roman. Methodology for teaching introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):123–128, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Ortega:1974:NAP

[Ort74] James M. Ortega. Numerical analysis position. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6 (3):86–87, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ottenstein:1976:AAD

[Ott76] Karl J. Ottenstein. An algorithmic approach to the detection and prevention of plagiarism. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(4):30–41, December 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/virtual-machines.bib.

Oliver:1973:SPS

[OWE73] Richard C. Oliver, H. E. Wright, and W. C. Elrod. Statistical parameters in a shear flow field. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):121– 124, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Pagan:1978:RCP

[Pag78] Frank G. Pagan. The role of a comparative programming languages course in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):22–23, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Palay:1976:SUT

[Pal76] Roger M. Palay. The structure and use of a test generating system designed to facilitate individually paced instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):100–103, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE– SIGCUE joint symposium on Computer science education.

Parnas:1972:CSE

[Par72] D. L. Parnas. A course on software engineering techniques. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):154–159, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pitts:1973:TYC

[PB73] Gerald N. Pitts and Barry Bateman. A two year computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):38–40, February 1973. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Pitts:1974:SOC

[PB74] Gerald N. Pitts and Barry Bateman. A software oriented computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):33–36, February 1974. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Pitts:1977:SAP

[PB77] Gerald N. Pitts and Barry L. Bateman. A structured approach to problem solving at the introduction level in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):168–172, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Platt:1976:EPC

[PC76] Joan Platt and Jeanne Curran. An experimental program in computer usage for secondary students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):84–85, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Powell:1979:CPE

[PDS^{+}79] James D. Powell, J. Richard Dennis, William Schaeffer, James E. Gray, and Robert Taylor. Computers in precollege education (panel discussion): What are colleges of education doing? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):167, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Pitts:1973:CSP

[PE73] Gerald N. Pitts and Roy S. Ellzey. Computer science — a professional degree. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(4):8–11, December 1973. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Peck:1975:PCS

[Pec75] Josephine Peck. A proposed Computer Science curriculum for twoyear colleges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):148–156, February 1975. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Peck:1977:USD

[Pec77] John C. Peck. The university — a systems development center for state government or how to solve the education vs. training problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):1–5, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Peelle:1976:LMR

[Pee76] Howard A. Peelle. Learning mathematics with recursive computer programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):116–130, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Phister:1976:PCD

[Phi76] M. Phister, Jr. A proposed course on data processing economics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):260–265, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Poore:1974:TCH

[PK74] Jesse H. Poore, Jr. and Carl J. Kropff. Teaching computer hardware organization with a simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(4):36–48, December 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Powell:1978:SVP

[PL78] J. D. Powell and M. J. Lee. Simulation as a vehicle for project experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):63–66, August 1978. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Poirot:1976:CDT

[Poi76] J. L. Poirot. A course description for teacher education in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):39–48, February 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Poirot:1978:CLE

[Poi78] J. L. Poirot. Computers in law enforcement: a service course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):32–34, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Poirot:1979:CES

[Poi79] James L. Poirot. Computer education in the secondary school: Problems and solutions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):101–104, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Pollack:1972:LSS

[Pol72] Seymour M. Pollack. A look at the second SIGCSE technical symposium. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(2):14–15, July 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pollack:1978:DUI

[Pol78] Seymour Pollack. Design and use of instructional tools for computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):1–6, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Posch:1979:MHS

[Pos79] Reinhard Posch. Modelling a hardware structure for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(2):60–68, June 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Powell:1978:UMC

[Pow78a] James D. Powell. Use of model curricula in reviewing established programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):12–15, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Powers:1978:DPP

[Pow78b] Michael J. Powers. The directed project: a poor man’s internship. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):97, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/ CSA Technical Symposium on Computer Science Education.

Powell:1979:AOI

[Pow79] James D. Powell. An application oriented introductory computer science sequence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):49–53, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Pavlak:1977:AFE

[PP77] Raymond Pavlak, Jr. and Adir Pridor. Assembler in a Fortran environment with a new debugging aid. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):63– 69, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prather:1976:ALD

[Pra76] Ronald E. Prather. Another look at the discrete structures course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):247–252, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Perry:1978:PEU

[PS78a] James M. Perry and Norman E. Sondak. The project experience in undergraduate computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10 (2):21–30, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pinkert:1978:OBD

[PS78b] J. Pinkert and K. Schenk. Only a Baker’s dozen. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):81–84, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Prather:1978:LLA

[PS78c] Ronald E. Prather and Judith D. Schlesinger. A lecture/laboratory approach to the first course in programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):115–118, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Pumplin:1975:SET

[Pum75] Bruce A. Pumplin. A simple example of the top-down approach to algorithm development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):35–37, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Purdy:1975:GCD

[Pur75] J. Gerry Purdy. A graduate course in database management. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):21–26, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Plum:1974:TSP

[PW74] T. W. S. Plum and G. M. Weinberg. Teaching structured programming attitudes, even in APL, by example. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6 (1):133–143, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Perry:1975:MCP

[PW75a] James T. Perry and Terry E. Weymouth. A modified chief programmer team approach to an operating systems class project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):31–39, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Porter:1975:PTA

[PW75b] Charles H. Porter and Nesa Labbe Wu. Programming for terminal applications. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):77–82, September 1975. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Radue:1976:TEF

[Rad76] J. E. Radue. On the teaching and evaluation of a Fortran service course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(2):32–35, June 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ragsdale:1976:MDP

[Rag76] Ronald G. Ragsdale. Multi-disciplinary programming exercises. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):295–297, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Ragsdale:1979:BPP

[Rag79] Ronald G. Ragsdale. A Basic program package for introducing the top-down approach to computer programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):113–117, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Ralston:1971:FFC

[Ral71] Anthony Ralston. Fortran and the first course in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(4):24–29, December 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Randolph:1975:CFS

[Ran75] Robert H. Randolph. Computer facility: Starter kit. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):15–19, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Rink:1977:FNU

[RD77] R. A. Rink and G. Droucas. A further note on using GPSS to study a student-operating system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):45–52, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reisman:1971:HCC

[Rei71] Sorel Reisman. A hypothetical computer configuration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(4):19–21, December 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reilly:1972:PEP

[Rei72] Edwin D. Reilly. Problems, exams, projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(2):28, July 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reisman:1973:SPP

[Rei73] Sorel Reisman. A survey of pedagogical programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(2):13–20, June 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ripley:1978:LCS

[RGH78] Jerald L. Ripley, Jarrell C. Grout, and Denis M. Hyams. The laboratory in a computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(2):12– 16, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rahimi:1973:ECS

[RH73] M. A. Rahimi and H. G. Hedges. Evolution of a computer science academic program in a college of engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5 (1):110–114, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Rheinboldt:1973:CRC

[Rhe73] Werner C. Rheinboldt. The CUPM recommendations on a computational mathematics program and on the impact of computers on undergraduate mathematics courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):83–85, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Rickman:1975:FSP

[Ric75] Jon Rickman. File structures position paper. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7 (2):48–50, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rine:1978:CCS

[Rin78] David C. Rine. Curricula in computer science and engineering: needs and alternatives. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):99–106, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Ripley:1975:CEP

[Rip75] G. David Ripley. A course in effective programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):102–108, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Rose:1971:ATA

[RJB71] Lawrence L. Rose, Neil D. Jones, and Bruce H. Barnes. Automata: a teaching aid for mathematical machines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(1):12– 20, March 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robertson:1975:PFC

[Rob75] Edward L. Robertson. The problems facing computer science education in developing nations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):56–60, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Rogers:1978:PTC

[Rog78] James L. Rogers. Problems of teaching computer history in introductory courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):239–242, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Rood:1972:PEP

[Roo72] Robert E. Rood. Problems, exams, and projects: Computer systems U302. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(4):32–44, December 1972. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosin:1971:FNC

[Ros71] Robert F. Rosin. Fortran and the new culture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3 (1):10–11, March 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosenberg:1976:ICS

[Ros76] Ivan M. Rosenberg. Introductory computer science courses a modular design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):49–61, February 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Roth:1972:CSC

[Rot72] R. Waldo Roth. A computer science curriculum for a liberal arts college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):37, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roth:1975:SFT

[Rot75] R. Waldo Roth. Students and faculty training in systems analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):67–73, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roth:1978:BCD

[Rot78] R. Waldo Roth. A beginning curriculum in data processing: for Native American students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):137–148, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Ruby:1976:SCS

[Rub76] Douglas A. Ruby. A survey on computer science curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):313–323, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Ruschitzka:1977:OSI

[Rus77] Manfred Ruschitzka. An operating systems implementation project for an undergraduate course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):77–84, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Sterling:1978:CSE

[SAK^{+}78] Theodor Sterling, F. K. Allotey, Asad Khailany, Maria Lucia Blanck Lisboa, Ignacio Mijares, and Edward Robertson. Computer science education in developing countries. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):179, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Salton:1973:IPC

[Sal73] Gerard Salton. Introductory programming at cornell. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):18–20, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Salako:1976:FMI

[Sal76] Abimbola Salako. A functional model of instructional programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):393–399, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Salomon:1977:THC

[Sal77] David Salomon. Teaching hardware and computer design in a software oriented computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):49– 50, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Sammet:1971:PTU

[Sam71] Jean E. Sammet. Proposed technological updating course for programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(1):25–33, March 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sampson:1974:IAS

[Sam74] Jeffrey R. Sampson. An introductory adaptive systems course for undergraduate computer science majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):148– 151, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Sammet:1975:DCG

[Sam75] Jean E. Sammet. Description of course given on “computers in the humanities and social science”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):133– 142, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Sangare:1978:CLI

[San78] Demba Sangare. Communication sur “l’enseignement de l’informatique dans les pays en voie de dévelopement”. (French) [Communication on “the teaching of computer science in developing countries”]. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):149–152, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Sapega:1975:DIC

[Sap75] A. E. Sapega. Development and implementation of a computing major program based on suggested ACM curricula at an undergraduate liberal arts college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):65–67, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Steinhorst:1974:CDU

[SB74] G. Cort Steinhorst and Barry L. Bateman. Curriculum development for an undergraduate course in data and information structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(2):20–23, June 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Smith:1976:GSU

[SB76] Robert L. Smith and Lee H. Blaine. A generalized system for university mathematics instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):280– 288, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Sanders:1976:SSC

[SBS76] William R. Sanders, Gerard V. Benbassat, and Robert L. Smith. Speech synthesis for computer assisted instruction: The MISS system and its applications. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):200–211, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Schwenkel:1971:RND

[Sch71] Frieder Schwenkel. Remarks on the Notre Dame computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(1):36–40, March 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schweppe:1973:DIM

[Sch73] Earl J. Schweppe. Dynamic instructional models of computer organizations and programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):26– 31, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Schulman:1977:TUC

[Sch77] Edward L. Schulman. Turning on the undergraduate computer science student: a RE-IPL suggestion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):178– 179, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Schneider:1978:IPC

[Sch78a] G. Michael Schneider. The introductory programming course in computer science: ten principles. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):107–114, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Schrage:1978:INC

[Sch78b] John F. Schrage. Industrial needs culminated in business applications course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):85, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Schneider:1979:YAS

[Sch79] G. Michael Schneider. Yet another survey of graduate programs in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(2):52–59, June 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schneider:1978:SSI

[SCM78] M. L. Schneider, N. S. Coulter, and J. C. Munson. A successful student internship program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):98, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Sebaugh:1976:SAI

[Seb76] Jeanne L. Sebaugh. The stepwise approach to introductory programming projects with examples. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):372– 381, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Semple:1973:ECS

[Sem73] Wolsey A. Semple. Evolution of a computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):115–118, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Senn:1974:POP

[Sen74] James A. Senn. A problem oriented pedagogy for computer language instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(4):26–29, December 1974. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Spence:1978:SAD

[SG78] J. Wayne Spence and Jarrell C. Grout. Systems analysis and design in a computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4): 24–27, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shaw:1972:ICC

[Sha72] Mary Shaw. Immigration course in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(2):16–27, July 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shaw:1973:ICC

[Sha73] Mary Shaw. Immigration course in computer science: teaching materials and 1972 schedule. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(2):26–32, June 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Swanson:1979:CDI

[SHLS79] Trevor Swanson, Richard Hatch, Lee Lane, and Norman Sondak. Curriculum development in information systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11 (1):202–206, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Shneiderman:1971:CSE

[Shn71] Ben Shneiderman. Computer science education and social relevance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(1):21–24, March 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shneiderman:1977:EIP

[Shn77] Ben Shneiderman. Evaluating introductory programming textbooks: a guide for students, instructors, authors and publishers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):56–58, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Shub:1979:IRS

[Shu79] Charles M. Shub. Isolated rural schools can have computers and teachers who can utilize them. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):105–112, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Senn:1979:BER

[SI79] James A. Senn and Blake Ives. Behavioral education requisites for application-oriented computer scientists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):195–201, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Sjoerdsma:1976:IPA

[Sjo76] Ted Sjoerdsma. An interactive pseudo-assembler for introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):342–349, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Sammet:1974:RFD

[SLC^{+}74] Jean E. Sammet, Robert Lechner, Richard G. Canning, Joseph Cunningham, George G. Dodd, Alex A. J. Hoffman, Anthony Ralston, and Frederic G. Withington. Recommended future directions for ACM. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(4):5–12, December 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sloan:1974:CAU

[Slo74] M. E. Sloan. Computer architecture in U.S. and Canadian electrical engineering departments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):111–115, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Slonneger:1976:CA

[Slo76] K. Slonneger. Computer attitudes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(4):26–29, December 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shapiro:1976:NAT

[SM76] Henry D. Shapiro and M. Dennis Mickunas. A new approach to teaching a first course in compiler construction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):158–166, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE– SIGCUE joint symposium on Computer science education.

Smith:1976:UDF

[Smi76] James A. Smith. The use of data flow diagrams in an introductory course in assembly language programming and machine architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(4):42–52, December 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Smillie:1978:ICS

[Smi78] K. W. Smillie. An introductory computing science course for liberal arts students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(2):63–65, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Smith:1979:SLA

[Smi79] James Smith. The small liberal arts college: a challenge for computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):220–223, February 1979. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Smoliar:1976:MS

[Smo76a] Stephen W. Smoliar. Meditations on style. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):75– 77, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Smoliar:1976:WFY

[Smo76b] Stephen W. Smoliar. What is a first-year computer science graduate student? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):304–306, February 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Snyder:1970:CSI

[Sny70] Mitchell Snyder. Computer science in Israel’s institutions of higher learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):69–74, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sobel:1976:HMC

[Sob76] Helen L. Sobel. How much CAI is hidden in commercial software? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):253–259, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Solomon:1976:LCC

[Sol76] Cynthia J. Solomon. Leading a child to a computer culture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):79–83, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Solntseff:1978:PLI

[Sol78] N. Solntseff. Programming languages for introductory computing courses: a position paper. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):119–124, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Solntseff:1979:CSM

[Sol79] N. Solntseff. A computer science microcomputer laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):149–154, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Sommerville:1977:MLP

[Som77] Ian Sommerville. Machine language programming in an undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):104– 107, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Sondak:1975:CSM

[Son75] Norman E. Sondak. A computer science minor for engineering and science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(4):19–24, December 1975. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sondak:1978:CSE

[Son78] Norman Sondak. A computer science educational software exchange. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):74–78, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Sterling:1970:EUI

[SP70] T. Sterling and S. Pollack. Experience with a “universal” introductory course in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):106– 112, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sterling:1974:TSI

[SP74] T. D. Sterling and S. V. Pollack. Teaching simulators or ideal teaching machines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(2):45–56, June 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Smith:1979:UFM

[SP79] Donald R. Smith and James T. Palmer. Universal fixed messages and the Rivest–Shamir–Adleman cryptosystem. Mathematika, 26 (1):44–52, 1979. CODEN MTKAAB. ISSN 0025-5793.

Stubbs:1973:IGE

[SR73] D. F. Stubbs and D. W. Roberts. Implementing Gauss elimination. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(3):18–24, September 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Smith:1976:SLP

[SR76] Carol Smith and Jon Rickman. Selecting languages for pedagogical tools in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):39– 47, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Sistare:1974:IDC

[SS74] John H. Sistare and Norman E. Sondak. Introduction to digital computer programming an IPI approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):184– 194, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Svoboda:1979:CAD

[SS79] Rudy Svoboda and Robert Sedlmeyer. Computer-aided design of graphical review lessons for audio-tutorial college algebra and trigonometry. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):232–235, February 1979. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Stoddard:1979:BDI

[SSL79] S. D. Stoddard, R. L. Sedlmeyer, and R. G. Lee. Breadth or depth in introductory computer courses: a controlled experiment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):41–44, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Stark:1972:CSN

[Sta72] Richard H. Stark. Computer science needs its laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):46–48, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Steinberg:1972:SPP

[Ste72] David I. Steinberg. A student project on political redistricting by computer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):86–89, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stern:1978:DBC

[Ste78] Myles Stern. A data base course for business students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):70–73, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/ CSA Technical Symposium on Computer Science Education.

Sticha:1973:ECS

[Sti73] Philip E. Sticha. Experience with computer science at a two year college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):41–44, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Stokes:1974:CSD

[Sto74a] Gordon E. Stokes. Computer science department service courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):28–29, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stokes:1974:SCP

[Sto74b] Gordon E. Stokes. Service course position paper. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6 (3):18–20, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stone:1975:ICS

[Sto75] Fred W. Stone. Innovative computer services at a developing institution. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):120–122, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Stoddard:1977:MSI

[Sto77] Spotswood D. Stoddard. Minicomputer selection and implementation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):6–12, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Stoddard:1978:MCC

[Sto78] Spotswood D. Stoddard. A mini computer course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):157–159, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Stubbs:1971:BDP

[SW71] D. F. Stubbs and N. W. Webre. Bachelor’s degree program in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(3):10–14, September 1971. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shapiro:1974:IVS

[SW74a] Stuart C. Shapiro and Douglas P. Witmer. Interactive visual simulators for beginning programming students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):11–14, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Sterlling:1974:SFO

[SW74b] T. D. Sterlling and J. J. Weinkam. The Simon Fraser one track universal curriculum for computing science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):28– 32, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Szalajka:1979:ITP

[SW79] Walter S. Szalajka and Philip Walch. Integrated theory and practice — an approach to the first computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):45–48, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Swigger:1976:LTP

[Swi76] Kathleen M. Swigger. Learning theory and political socialization: a module for computer assisted instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):289– 294, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Taylor:1977:TPB

[Tay77] Robert P. Taylor. Teaching programming to beginners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):88–92, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Taylor:1979:GRT

[Tay79] Robert P. Taylor. Graduate remedial training in computing for educators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):168–173, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Tam:1977:PET

[TB77] W. C. Tam and S. N. Busenberg. Practical experience in top-down structured software production in an academic setting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):31–36, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Tcheng:1975:UCS

[Tch75] Mike T. Tcheng. The undevelopment of a computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):74–77, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tharp:1979:ICP

[Tha79] Alan L. Tharp. Interactive computing in a project-oriented file organization course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):82–86, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Thomas:1974:CAC

[Tho74] Richard T. Thomas. Computer architecture in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):116–120, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Thorhallsson:1975:HCA

[Tho75] Jon Thorhallsson. High-school curriculum adopts first semester college computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):17–18, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:1979:DIC

[Tho79] Paul A. V. Thomas. DRAW(IT)N — a computer graphics education package. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):228–231, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Thurber:1979:CSC

[Thu79] Kenneth J. Thurber. A course structure: computer communications subsystems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(4):15–24, December 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Trickey:1979:EOM

[TL79] H. Trickey and E. S. Lee. Experience with OLGA, a maximal language for the minimal PDP-11 minicomputer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11 (1):240–244, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Tremblay:1974:FCD

[TM74] J. P. Tremblay and R. Manohar. A first course in discrete structures with applications to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):155–160, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Thompson:1977:ILS

[TMWD77] William B. Thompson, Thomas G. Moher, William T. Wood, and David W. DeHerder. An instrumental laboratory for systems software development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):38–40, December 1977. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tartar:1972:UEC

[TP72] J. Tartar and J. P. Penny. Undergraduate education in computing science: some immediate problems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):1–7, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tracz:1974:UAP

[Tra74] Will Tracz. The use of ATOPSS for presenting elementary operating system concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):74–78, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Traub:1976:SGO

[Tra76] J. F. Traub. Some general observations on Ph.D. production in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(4):8–9, December 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Trombetta:1979:TPA

[Tro79] Michael Trombetta. On testing programming ability. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(4):56–60, December 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tremblay:1975:ICD

[TS75] J. P. Tremblay and P. G. Sorenson. An introductory course in data structures with applications. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):50–57, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thalmann:1978:DCB

[TT78a] Nadia Thalmann and Daniel Thalmann. Direct connection between Compiling Techniques and Databases courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):131–135, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Thalmann:1978:UPT

[TT78b] Nadia Thalmann and Daniel Thalmann. The use of PASCAL as a teaching tool in introductory, intermediate and advanced computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):277–281, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Unger:1976:IAC

[UA76] E. A. Unger and N. Ahmed. An instructionally acceptable cost effective approach to a general introductory course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(2):28–31, June 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Varanelli:1978:EIU

[Var78] Andrew Varanelli. An effectiveness index for use in university computing center management. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):232–238, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Varanelli:1979:CEM

[Var79a] Andrew Varanelli, Jr. Computer education in the management curriculum at Pace University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):2–5, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Varden:1979:TPD

[Var79b] Stuart A. Varden. Teaching program design through program structure maps. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):182–186, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Vawter:1976:UIT

[Vaw76] Richard Vawter. The use of information theory and personal probabilities in computer based learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):284– 288, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

vanDam:1976:SPA

[vDDDN76] Andries van Dam, Jens M. Dill, Douglas F. Dixon, and David S. Notkin. Structured programming in assembly language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(4):53–67, December 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

vandeRiet:1979:LLC

[vdR79] R. P. van de Riet. The little language compiler course project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(4):31–48, December 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

vanDam:1974:SIA

[vDSMM74] Andries van Dam, Charles M. Strauss, Clement McGowan, and Jean Morse. A survey of introductory and advanced programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):174–183, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Vickers:1971:DCS

[Vic71a] F. D. Vickers. Data on computer science departments/curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(4):40–45, December 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vickers:1971:ISL

[Vic71b] F. D. Vickers. An information system for a large course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(3):7–10, September 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vishnubhotla:1977:LAT

[Vis77] Sarma R. Vishnubhotla. A logical approach to teach digital computer design at logic and systems level. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):13– 21, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Vishnubhotla:1978:PTM

[Vis78] Sarma R. Vishnubhotla. A project to teach microprogrammed asynchronous system design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):291–296, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Weiner:1974:AMI

[WA74] Leonard H. Weiner and William P. Allaire. The author-mode interface for the ASAG assignment generating system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):15–20, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Wagener:1975:SFP

[Wag75] J. L. Wagener. Structured FORTRAN programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):206–211, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Wainwright:1979:SFC

[Wai79] Roger Wainwright. A survey of faculty computer experience, usage, needs, literacy and attitudes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(2):27–35, June 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Waksman:1978:IRG

[Wak78] Abraham Waksman. On the introduction of report generating system design concepts in an introductory course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10 (2):42–43, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1973:CSC

[Wal73] Terry M. Walker. Computer science curricula survey. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(4):19–28, December 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1974:ECS

[Wal74] Terry M. Walker. An evaluation of computer science curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(3):81–85, September 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ward:1975:CDE

[War75a] Darrell L. Ward. A cooperative development effort of a computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):78–83, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wardle:1975:CSP

[War75b] Caroline Wardle. A computer science program at a college with limited resources. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):68–70, February 1975. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Waxman:1975:RBD

[Wax75] Jerry Waxman. Reflections on B3, discrete structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):51–54, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). See reply [Kor75].

Weaver:1978:MCS

[Wea78] Alfred C. Weaver. Microcomputers in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):171–176, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Weaver:1979:UMT

[Wea79] Alfred C. Weaver. Using microcomputers to teach computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):141–143, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Webb:1973:STP

[Web73] G. L. Webb. Some thoughts on program structure. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(3):8–12, September 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Webre:1975:MPB

[Web75] Neil Webre. The Master’s practicum: a bridge over the Industry/ University Gap. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):79–82, February 1975. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Wegner:1970:STG

[Weg70] Peter Wegner. Some thoughts on graduate education in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(4):30–40, September/October 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wegner:1971:PCS

[Weg71] Peter Wegner. Problem of computer science education in small colleges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(3):15–18, September 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weiner:1973:MGA

[Wei73] Leonard H. Weiner. Machine generation of assignments for a mass education introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):181–185, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Weiner:1977:FCS

[Wei77] Leonard H. Weiner. Foundations of computer science for entrylevel graduate students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):6–9, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Weiner:1978:RSP

[Wei78] Leonard H. Weiner. The roots of structured programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):243–254, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Welch:1976:TPD

[Wel76] John T. Welch, Jr. Tab, a program design language for the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):48–53, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

West:1972:PAI

[Wes72] Sheridan J. West. Providing adequate instruction to data processing students spread over a wide geographical area. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):24–26, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1973:PPO

[WH73] Justin C. Walker and Charles E. Hughes. POPSS — a parametric operating system simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):166–169, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Whithed:1972:TAD

[Whi72] Marshall H. Whithed. Technological analysis and democratic policy-making. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):75–85, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Whitlock:1977:DID

[Whi77] Lawrence R. Whitlock. The design and implementation of a dynamically tailored examination. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):59–62, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Willoughby:1972:UPD

[Wil72] Theodore C. Willoughby. An upside down PH.D. program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(4):8–10, December 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Willoughby:1973:SAT

[Wil73] Theodore C. Willoughby. Student attitudes toward computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):145–148, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Williams:1976:ECA

[Wil76a] Kenneth Williams. An experimental course in advanced programming methods. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(4):15–18, December 1976. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Willis:1976:CSC

[Wil76b] Neil Willis. Computing science courses — training or education? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):298–303, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Williams:1978:SCT

[Wil78] Kenneth Williams. SIGCSE/CSA 1978 technical symposium report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(2):8–9, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Winslow:1977:SIC

[Win77] L. Winslow. A structured introductory computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):165–167, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Winner:1978:ICC

[Win78] Robert I. Winner. An intermediate course on computer systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):210–212, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Woolley:1974:LSL

[WM74] John D. Woolley and Leland R. Miller. LINUS: a structured language for instructional use. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):125–128, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Wolfe:1976:AQG

[Wol76a] John H. Wolfe. Automatic question generation from text — an aid to independent study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):104–112, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Wolkov:1976:TCB

[Wol76b] David Wolkov. Teaching a computer-based college level statistics course in a secondary school. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):131– 136, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Woodhouse:1978:PJP

[Woo78] D. Woodhouse. Programming the Josephus problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):56–58, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Worlana:1978:UAC

[Wor78] Peter B. Worlana. Using the ACM computer science curriculum recommendations in a liberal arts college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):16– 19, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weymouth:1977:MOS

[WP77] Terry E. Weymouth and James T. Perry. A multiprogramming operating system for a minicomputer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):69– 72, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Winkel:1978:CSP

[WP78a] David Winkel and Franklin Prosser. Computer science is the proper home for digital hardware design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1): 153–156, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Winrich:1978:PIC

[WP78b] Lonny B. Winrich and Charles G. Petersen. A pragmatic implementation of Curriculum ’77 at the University of Wisconsin: La Crosse. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(4):37–41, December 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wu:1974:CTM

[Wu74] Chih Wu. CAI tutorial method of teaching thermodynamics at U.S. Naval Academy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):97–100, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Weinberg:1974:SPC

[WW74] Bernhard Weinberg and Leonard H. Weiner. A systems programming course using the HMS 5050, a counterfeit, hands-on, largescale computer system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):64–73, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Webster:1973:CSO

[WZ73] Dennis B. Webster and Victor A. Zaloom. A computer science option in industrial engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):125– 127, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Yelowitz:1978:PAS

[Yel78] Lawrence Yelowitz. A project approach to structure and correctness in Pitt’s second computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10 (3):50–54, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Yeh:1973:NDT

[YGM73] Raymond T. Yeh, Donald I. Good, and David R. Musser. New directions in teaching the fundamentals of computer science — discrete structures and computational analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):60–67, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Zelkowitz:1972:SRC

[Zel72] Marvin Zelkowitz. Space requirements for computer programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(2):10–11, July 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zimmerman:1971:CFE

[Zim71] E. C. Zimmerman. Computing facilities and efforts at a small liberal arts college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(1):34–35, March 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zinn:1970:ASS

[Zin70] Karl Zinn. Abstracts of SJCC ’70 session on computers in education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(1):8–10, March 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).